### v2

# On the structures of a monoid of triangular vector- permutation polynomials, its group of units and its induced group of permutations

Abstract
Let $n>1$ and let $R$ be a commutative ring with identity $1\neq 0$ and $R[x_{1},\ldots,x_{n}]^{n}$ the set of all $n$-tuples of polynomials of the form $(f_{1},\ldots,f_{n}),$ where $f_{1},\ldots,f_{n}\in R[x_{1},\ldots,x_{n}]$. We call these $n$-tuples vector-polynomials. We define composition on $R[x_{1},\ldots,x_{n}]^{n}$ by
$\vec{g}\circ\vec{f}=(g_{1}(f_{1},\ldots,f_{n}),\ldots,g_{n}(f_{1},\ldots,f_{n})),\text{ where }\vec{f}=(f_{1},\ldots,f_{n}),\vec{g}=(g_{1},\ldots,g_{n}).$
In this paper, we investigate vector-polynomials of the form
$\vec{f}=(f_{0},f_{1}+x_{2}g_{1},\ldots,f_{n-1}+x_{n}g_{n-1}),$
where $f_{0}\in R[x_{1}]$ permutes the elements of $R$ and $f_{i},g_{i}\in R[x_{1},\ldots,x_{i}]$ such that each $g_{i}$ maps $R^{i}$ into the units of $R$ ($i=1,\ldots,n-1$). We show that each such vector-polynomial permutes the elements of $R^{n}$ and that the set of all such vector-polynomials $\mathcal{MT}_{n}$ is a monoid with respect to composition. We also show that $\vec{f}$ is invertible in $\mathcal{MT}_{n}$ if and only if $f_{0}$ is an $R$-automorphism of $R[x_{1}]$ and $g_{i}$ is invertible in $R[x_{1},\ldots,x_{i}]$ for $i=1,\ldots,n-1$. When $R$ is finite, the monoid $\mathcal{MT}_{n}$ induces a finite group of permutations of $R^{n}$. Moreover, we decompose the monoid $\mathcal{MT}_{n}$ into an iterated semi-direct product of $n$ monoids. Such a decomposition allows us to obtain similar decompositions of its group of units and, when $R$ is finite, of its induced group of permutations. Furthermore, the decomposition of the induced group helps us to characterize some of its properties.
1. Introduction
Let $R$ be a commutative ring (not necessarily finite) with unity and let $R[x_{1},\ldots,x_{n}]$ be its polynomial ring in $n$ variables. A vector-polynomial $\vec{f}$ is an $n$-tuple $(f_{1},\ldots,f_{n})$ of polynomials over $R$ in the variables $x_{1},\ldots,x_{n}$, i.e., $\vec{f}=(f_{1},\ldots,f_{n})$ with $f_{i}\in R[x_{1},\ldots,x_{n}]$ for $i=1,\ldots,n$. It is well- known that every vector-polynomial $\vec{f}$ defines an $R$-endomorphism $\phi_{\vec{f}}$ of $R[x_{1},\ldots,x_{n}]$ by $\phi_{\vec{f}}(g)=g(f_{1},\ldots,f_{n})$ for any $g\in R[x_{1},$$,x_{n}]$. Furthermore, $\vec{f}$ is an $R$-automorphism of $R[x_{1},$$,x_{n}]$ if and only if it is invertible with respect to composition. The set of all these automorphisms is a group, which is denoted by $GA_{n}$. Vector-polynomials are commonly referred to as polynomial maps in the literature. We are interested not only in the vector-polynomial (polynomial map) but also in its induced function (map) on $R^{n}$ via evaluation. This distinction helps avoid confusion between an $n$-tuple of polynomials and a single polynomial in this context.
Several results have been achieved concerning these $R$-endomorphisms of $R[x_{1},\ldots,x_{n}]$, mainly considering the group $GA_{n}$, its subgroups, and its elements. The developed ideas cover several aspects of investigations, including the Jacobian Conjecture and equivalent statements (see for example and the references therein [24]), the inclusion of the tame group (see for example [11, 26, 22]), and the study of structures of some subgroups of $GA_{n}$ (see [2, 28]).
Previous works have primarily focused on the case where $R$ is an infinite field with characteristic zero. Nevertheless, researchers have also considered finite fields to investigate similar questions (see [14, 16, 12]), disproving some famous results known to be true in the infinite case (see [9, 16]), and reformulating some others (for instance [15]). Notably, the permutations induced by the elements of $GA_{n}$ on $R^{n}$ have a powerful impact.
In this paper, we explore the structure of a monoid of triangular vector-polynomials with a specific property. We consider the monoid of all triangular vector-polynomials of the form $(f_{0},f_{1}+x_{2}g_{1},\ldots,f_{n-1}+x_{n}g_{n-1})$ with the condition that $f_{0}$ is a permutation polynomial on $R$ and $g_{i}$ is a unit-valued on $R$ for $i=1,\ldots,n-1$. The elements of this monoid define permutations of $R^{n}$. Additionally, we observe that the group of units of this monoid generally contains properly the (lower) Jonquières (classical) triangular group unless $R$ is a reduced ring. In a reduced ring the invertible polynomials coincide with the units of the ring $R$, and in this case, the group of the triangular monoid is just the Jonquières group (see Remark 4.11).
The monoid under consideration can be decomposed into an iterated semi-direct product of $n$ monoids by iteration (see Theorem 4.5). Surprisingly, this decomposition admits a similar decomposition of the group of units of the triangular monoid into an iterated semi-direct product of $n$ groups by finding the units of those factor monoids in the first decomposition (Theorem 4.10). Further, when $R$ is finite, the triangular monoid induces a group of permutations on $R^{n}$, which is its homomorphic image via the natural epimorphism $\pi_{n}$ (defined in Remark 2.12). Again, we obtain an iterated semi-direct product decomposition of the induced group just by taking images (via $\pi_{n}$) of the factor monoids in that decomposition of the triangular monoid (Theorem 4.15). Such decomposition allows us to characterize the solvability and the nilpotency of this group restricted (by the Chinese Reminder Theorem) to the local case (Theorem 4.21).
It is remarkable that in the case where $R$ is an algebraically closed field of characteristic zero, the triangular monoid, its group of units, and its induced group all coincide with the Jonquières triangular group. This is because unit-valued polynomials are just the non-zero elements of $R$, and there is no non-linear permutation polynomial on $R$ (see Proposition 4.12). This motivates us to suggest a modification for the tameness problem in the general case by replacing the Jonquières with our triangular monoid, and then posing some related questions (see Section 6).
Having previously worked on polynomial functions and polynomial permutations over finite commutative rings (in one variable) rather than polynomial mappings (in several variables), we can ensure that our motivation to study the triangular monoid emanates from a simple question on polynomial permutations. In other words, given a finite ring $R$ and $R_{n}$ is its dual numbers extension in $n$ variables. Then, does every polynomial permutation on $R_{n}$ can be viewed as a vector-permutation polynomial on the base ring $R$? More generally, can the group of permutation polynomials on $R_{n}$ be viewed as a group of permutations of $R^{n+1}$ induced vector-polynomials in $n+1$ variables? We answer this question affirmatively and provide more details (see Section 5).
The structure of our paper is as follows: in Section 2, we introduce our notation and basic definitions, and we recall some facts about polynomials and polynomial functions. In section 3, we describe the elements of the triangular monoid and its units and also obtain some simple relations for them. Section 4 consists of three subsections, the first subsection contains the decomposition of the triangular monoid into an iterated semi-direct product of monoids, while the other two subsections are devoted to the decompositions of its group of units and its induced group of permutations, respectively. In section 5, we consider the problem of embedding the monoid of permutation polynomial (in one variable) of certain ring extensions of the ring $R$ into the triangular monoid. Finally, in Section 6, we formulate the tameness problem and suggest some related questions.
2. Preliminaries
Throughout this paper, let $R$ be a commutative ring with $1\neq 0$, $R^{\times}$ its group of units, and whenever $R$ is a finite local ring, $M$ will denote its unique maximal ideal. Also, we use $\mathbb{F}q$ to denote its residue field $R/M$ or, more generally, a finite field of $q$ elements.
Let $n\geq 1$. By $n$, we denote the number of variables $x_{1},\ldots,x_{n}$; and to distinguish the elements of $R[x_{1},\ldots,x_{n}]^{n}$ from those of $R[x_{1},\ldots,x_{n}]$, we use the symbol $\vec{f}$ to indicate an $n$-tuple $(f_{1},\ldots,f_{n})$ of polynomials $f_{1},\ldots,f_{n}\in R[x_{1},\ldots,x_{n}]$. In a similar manner we use symbols like $\vec{F}$ to indicate function $\vec{F}\colon R^{n}\longrightarrow R^{n}$, or equivalently to indicate an $n$-tuple $(F_{1},\ldots,F_{n})$ of functions $F_{1},\ldots,F_{n}\colon R^{n}\longrightarrow R$, i.e. $\vec{F}=(F_{1},\ldots,F_{n})$. It is well-known that a vector-polynomial $\vec{f}=(f_{1},\ldots,f_{n})$ defines an $R$-endomorphism of $R[x_{1},\ldots,x_{n}]$: $\phi_{\vec{f}}\colon R[x_{1},\ldots,x_{n}]\longrightarrow R[x_{1},\ldots,x_{n}]$, $\phi_{\vec{f}}(h)=h(f_{1},\ldots,f_{n})$. Such an endomorphism is called an $R$-automorphism if there is a vector-polynomial $\vec{g}=(g_{1},\ldots,g_{n})$ such that
$\vec{f}\circ\vec{g}=(g_{1}(f_{1},\ldots,f_{n}),\ldots,g_{n}(f_{1},\ldots,f_{n}))=(x_{1},\ldots,x_{n}).$
Here, we notice that the above operation is anti the normal composition. However, we will adapt later the normal one (see Definition 2.6). The reason is that we are also interested in the (vector)-functions induced by (vector)-polynomials, and we require that the operation on polynomials and their functions be consistent. Also, it is noticeable here that despite any one of these operations being under consideration in the monoid ${R[x_{1},\ldots,x_{n}]}^{n}$, the terms invertible, left invertible and right invertible are equivalent [24].
Furthermore, we note that if $\vec{F}$ is vector-function obtained from the $R$-automorphism $\vec{f}$ via evaluation, that is,
$\vec{F}(a,\ldots,a_{n})=\vec{f}(a_{1},\ldots,a_{n})=(f_{1}(a_{1},\ldots,a_{n}),\ldots,f_{n}(a_{1},\ldots,a_{n})),\text{ for every }(a_{1},\ldots,a_{n})\in R^{n},$
permutes the elements of $R^{n}$. However, the converse is not true. For example, consider the case $n=1$, $R=\mathbb{F}_{q}$, and $\vec{f}(x)=f(x)=x^{q}$. Evidently, $\vec{F}=F=id_{R^{n}}$; but, there is no polynomial $g$ such that $f\circ g=x$.
Notation 2.1.
For $k\geq 1$, let $R_{[k]}$ denote the polynomial ring $R[x_{1},\ldots,x_{k}]$ and let $R_{[k]}^{\times}$ denote its group of units. But, when $k=1$, we use $R[x]$ (or $R[x_{1}]$) and $R[x]^{\times}$ (or $R[x_{1}]^{\times}$) to denote the polynomial ring in one variable and its group of units rather than writing $R_{[1]}$ and $R_{[1]}^{\times}$, respectively.
Definition 2.2.
Let $R$ be a commutative ring, and $f,g\in R_{[n]}$. Then:
  1. (1)
    The polynomial $f$ gives rise to a polynomial function $F\colon R^{n}\longrightarrow R$ by substitution for the variables, that is, $F(a_{1},\ldots,a_{n})=f(a_{1},\ldots,a_{n})$ for every $(a_{1},\ldots,a_{n})\in R^{n}$. Further, if $F$ maps $R$ into $R^{\times}$ the group of units of $R$, we call $f$ a unit-valued polynomial and $F$ is a unit-valued polynomial function (on $R$).
  2. (2)
    If $g(a_{1},\ldots,a_{n})=f(a_{1},\ldots,a_{n})$ for every $(a_{1},\ldots,a_{n})\in R^{n}$, this means that $f$ and $g$ induce the same function $F$ on $R$ and we write $f\operatorname{{}\triangleq{}}g$ on $R$. In this case, we say $f$ and $g$ are equivalent on $R$.
  3. (3)
    We denote by $\mathcal{F}(R^{n})$ the set of all polynomial functions (in $n$ variables) on $R$.
  4. (4)
    We denote by $UV(R_{[n]})$ the set of all unit-valued polynomials in $n$ variables.
From Definition2.2, we have easily the following fact.
Fact 2.3.
Let $R$ be a commutative ring. Then
  1. (1)
    $UV(R_{[n]})$ is a submonoid of $(R_{[n]},``\circ")$;
  2. (2)
    $\mathcal{F}(R^{n})$ is a monoid with respect to composition;
  3. (3)
    $\mathcal{F}(R^{n})$ is a commutative ring with pointwise operation. Furthermore, $\mathcal{F}(R^{n})$ is finite whenever $R$ is finite, and in this case the group of units of $\mathcal{F}(R^{n})^{\times}$ consists of all unit-valued polynomial functions.
Remark 2.4.
  1. (1)
    By Definition 2.2, every unit polynomial in a polynomial ring is a unit-valued polynomial but not vice versa. For example, consider the case $R$ is a finite local ring and $f(x)=(x^{q}-x)+1$.
  2. (2)
    When $R$ is a domain, the invertible polynomials are just the units of $R$.
  3. (3)
    When $R$ contains a non-zero nilpotent, it contains a nilpotent $r\neq 0$ such that $r^{2}=0$. But, then the polynomial $f(x)=rx+1$ admits $rx-1$ as multiplicative inverse.
  4. (4)
    Let $R$ be a commutative ring and $f(x)=\sum_{i=0}^{k}a_{i}x^{i}\in R[x]$. Then $f$ is a unit in $R[x]$ if and only if $a_{0}$ is a unit of $R$ and $a_{i}$ is nilpotent for $i=1,\ldots,k$ (see for example [3, Proposition. 1.3.1]).
Definition 2.5.
Let $R$ be a finite commutative ring. We call the group $\mathcal{F}(R^{n})^{\times}$ the group unit-valued polynomial functions (in $n$ variables) on $R$.
Definition 2.6.
Let
$\vec{f}=(f_{1},\ldots,f_{n}),\text{ and }\vec{g}=(g_{1},\ldots,g_{n}),\text{{ where} }f_{i},g_{i}\in R_{[n]}\text{ for }i=1,\ldots,n.$
Then we define $f_{i}(\vec{g})=f_{i}(g_{1},\ldots,g_{n})$ for $i=1,\ldots,n$; and
$\vec{f}\circ\vec{g}=(f_{1}(\vec{g}),\ldots,f_{n}(\vec{g}))=(f_{1}(g_{1},\ldots,g_{n}),\ldots,f_{n}(g_{1},\ldots,g_{n})).$
From the Definition 2.6, one easily prove the following fact.
Fact 2.7.
For $i=1,\ldots,n$, let $f_{i}\in R_{[i]}$. Then
  1. (1)
    $(f_{1},f_{2},\ldots,f_{n})=(f_{1},x_{2},\ldots,x_{n})\circ\cdots\circ(x_{1},x_{2},\ldots,f_{n})$;
  2. (2)
    $(x_{1},\ldots,x_{i},f_{i+1},\ldots f_{n})\circ\vec{g}=\vec{g}\circ(x_{1},\ldots,x_{i},f_{i+1}(\vec{g}),\ldots f_{n}(\vec{g}))$, where
    $\vec{g}=(x_{1},\ldots,x_{j-1},f_{j},x_{j+1},\ldots,x_{n})$ and $j\leq i$.
Definition 2.8.
Let $R$ be a commutative ring, and $f_{1},\ldots,f_{n};g_{1},\ldots,g_{n}\in R_{[n]}$. Then
  1. (1)
    The vector-polynomial $\vec{f}=(f_{1},\ldots,f_{n})$ gives rise to a vector-polynomial function of $R^{n}$, $\vec{F}\colon R^{n}\longrightarrow R^{n}$ by substitution for the variables, that is,
    $\vec{F}(a_{1},\ldots,a_{n})=\vec{f}(a_{1},\ldots,a_{n})=(f_{1}(a_{1},\ldots,a_{n}),\ldots,f_{n}(a_{1},\ldots,a_{n})),\text{ for every }(a_{1},\ldots,a_{n})\in R^{n}.$
    If $\vec{F}$ is bijective, we call $\vec{F}$ a vector-polynomial permutation, $\vec{f}$ a vector-permutation polynomial and $f_{i}$ a permutation polynomial for $i=1,\ldots,n$.
  2. (2)
    If $\vec{g}=(g_{1},\ldots,g_{n})$ and $\vec{g}(a_{1},\ldots,a_{n})=\vec{f}(a_{1},\ldots,a_{n})$ for every $(a_{1},\ldots,a_{n})\in R^{n}$, this means that $\vec{f}$ and $\vec{g}$ induce the same function $F$ on $R^{n}$ and we write $\vec{f}\operatorname{{}\triangleq{}}\vec{g}$ on $R^{n}$.
  3. (3)
    We denote by $\vec{\mathcal{F}}(R^{n})$ the set of all vector-polynomial functions of $R^{n}$.
Remark 2.9.
  1. (1)
    It is interesting to note that every invertible vector-polynomial is a vector-permutation polynomial and that the converse is not always true; however, in the case where $R$ is an algebraically closed field of characteristic zero, vector-permutation polynomials are just the invertible vector-polynomials (see [6, Theorem 2.2] and [24, Page. 80]).
  2. (2)
    When $n>1$, a permutation polynomial $f\in R_{[n]}$ can not alone induce a permutation of $R^{n}$ by Definition 2.8. However, when $n=1$, the concept of permutation polynomial and vector-permutation polynomial coincide. Permutation polynomials in one variable and their induced permutations play a vital rule in this context. For this reason we collect the related terminology in the following definition.
Definition 2.10.
Let $R$ be a commutative ring, and let $f\in R[x]$ and $F\colon R\longrightarrow R$ be its induced function. Then
  1. (1)
    $f$ is called a permutation polynomial whenever $F$ is a permutation of $R$. In this case, $F$ is called polynomial permutation.
  2. (2)
    By $\mathcal{MP}(R)$, we denote the monoid of permutation polynomials on $R$ consisting of all permutation polynomials.
  3. (3)
    The group of units of $\mathcal{MP}(R)$ consists of all $R$-automorphisms of $R[x]$, which we denote by $Aut_{R}(R[x])$.
  4. (4)
    If $R$ is finite, the set of all polynomial permutations on $R$ is a finite group which we denote by $\mathcal{P}(R)$.
Remark 2.11.
In [7], Gilmer showed that a polynomial $f\in R[x]$ is an $R$ automorphism of $R[x]$ if and only if $f(x)=a_{0}+a_{1}x+\cdots+a_{m}x^{m}$ with $a_{1}$ is unit and $a_{i}$ is nilpotent for every $i\geq 2$. It follows, when $R$ is reduced, i.e. $R$ has no non-trivial nilpotents elements, that $Aut_{R}(R[x])$ consists only of invertible linear polynomials.
Remark 2.12.
There is a natural epimorphism $\pi_{n}\colon R_{[n]}^{n}\longrightarrow\vec{\mathcal{F}}(R^{n})$ with respect to composition which maps every vector-polynomial into its induced vector-polynomial function $\vec{F}$.
Lemma 2.13.
Let $x_{1},\ldots,x_{k+1}$ be variables, and let $f,g,h,u\in R_{[k]}$ such that $h$ and $u$ are unit-valued polynomials. Then $f+x_{k+1}u\operatorname{{}\triangleq{}}g+x_{k+1}h$ on $R$ if and only if $f\operatorname{{}\triangleq{}}g$ and $u\operatorname{{}\triangleq{}}h$ on $R$.
Proof.
Suppose that $f+x_{k+1}u\operatorname{{}\triangleq{}}g+x_{k+1}h$ on $R$. Then setting $x_{k+1}=0$ yields that $f\operatorname{{}\triangleq{}}g$ on $R$. Hence, $x_{k+1}u\operatorname{{}\triangleq{}}x_{k+1}h$ on $R$. But, then $x_{k+1}=1$ implies that $u\operatorname{{}\triangleq{}}h$ on $R$. The other implication is obvious. ∎
Fact 2.14.
Let $R$ be a finite commutative ring. Let $\vec{F}$ be the vector-polynomial permutation of $R^{n}$ induced by the vector-polynomial $\vec{f}=(f_{1},\ldots,f_{n})$, where $f_{1},\ldots,f_{n}\in R_{[n]}$. If the inverse of $\vec{F}$, $\vec{F}^{-1}$, is induced by a vector-polynomial $\vec{g}=(g_{1},\ldots,g_{n})$, where $g_{1},\ldots,g_{n}\in R_{[n]}$, then $f_{i}\circ g_{i}\operatorname{{}\triangleq{}}x_{i}$ on $R$ for $i=1,\dots,n$.
In the following, we recall the definition of a short exact sequence of groups and its relation with the semi-direct product of groups. A finite sequence of groups and homomorphisms
[…]
is a short exact sequence if $\operatorname{Im}\alpha=\ker\beta$. Further, if $\alpha$ is the inclusion function, $i$, $B$ is an extension of $A$ by $C$. This allows us to give a definition of the semi-direct product of groups by means of a short exact sequence.
Definition 2.15.
[20, Page 760] An extension
[…]
is split if there exists a homomorphism $j\colon C\rightarrow B$ such that $pj=id_{C}$. In this case, the group $B$ is called the semi-direct product of $A$ by $C$, which is denoted by $A\rtimes C$.
3. The monoid of triangular vector-permutation polynomials
In this section, we give the construction of a monoid of triangular vector-polynomials. Recall from Definition 2.2 that $UV(R_{[k]})$ stands for the monoid of unit-valued polynomials in $k$ variables.
Lemma 3.1.
Let $R$ be a commutative ring. Let $\vec{f}=(f_{1},f_{2}+x_{2}g_{2},\ldots,f_{n}+x_{n}g_{n})$ and $\vec{h}=(h_{1},h_{2}+x_{2}w_{2},\ldots,h_{n}+x_{n}w_{n})$, where $f_{1},h_{1}\in\mathcal{MP}(R)$, $f_{i},h_{i}\in R_{[i-1]}$ and $u_{i},w_{i}\in UV(R_{[i-1]})$ with $i=2,\ldots,n-1$. Then $\vec{f}\circ\vec{h}=(l_{1},l_{2}+x_{2}r_{2},\ldots,l_{n}+x_{n}r_{n}),$ for some $l_{1}\in\mathcal{MP}(R)$, $l_{i}\in R_{[i-1]}$ and $r_{i}\in UV(R_{[i-1]})$ for $i=2,\ldots,n$.
Proof.
Let $\vec{h}\in\mathcal{MT}_{n}$ and $\vec{f}$ as above. Then, by Definition 2.6,
$\vec{f}\circ\vec{h}=(l_{1},l_{2}+x_{2}r_{2},\ldots,l_{n}+x_{n}r_{n}),$ (1)
where
$\begin{matrix}l_{1}(x_{1})=f_{1}(h_{1}(x_{1}))\\ l_{i}(x_{1},\ldots,x_{i-1})=f_{i}(\vec{v_{i}})+h_{i}(x_{1},\ldots,x_{i-1})g_{i}(\vec{v_{i}}),\\ r_{i}(x_{1},\ldots,x_{i-1})=w_{i}(x_{1},\ldots,x_{i-1})g(\vec{v_{i}})\text{ and }\\ \vec{v_{i}}(x_{1},\ldots,x_{i-1})=(h_{1}(x_{1}),h_{2}(x_{1})+x_{2}w_{1}(x_{1}),\ldots,h_{i-1}(x_{1},\ldots,x_{i-2})+x_{i-1}w_{i-1}(x_{1},\ldots,x_{i-2})),\\ i=2,\ldots,n.\end{matrix}$ (2)
But, then $l_{1}=f_{1}\circ h_{1}\in\mathcal{MP}(R)$ since $f_{1},h_{1}\in\mathcal{MP}(R)$; and for $i=2,\ldots,n$, $r_{i}=w_{i}g(\vec{v_{i}})\in UV(R_{[i-1]})$ since $w_{i},g(\vec{v_{i}})\in UV(R_{[i-1]})$. ∎
Recall from Remark 2.12 that $\pi_{n}$ is the epimorphism assigns to each vector-polynomial its induced function on $R^{n}$.
Lemma 3.2.
Let $R$ be a commutative ring. Let $f_{1}\in\mathcal{MP}(R)$ and for $i=2,\ldots,n$ let $f_{i}\in R_{[{i-1}]}$ and $g_{i}\in UV(R_{[{i-1}]]})$. Then the vector-polynomial: $\vec{f}=(f_{1},f_{2}+x_{2}g_{2},\ldots,f_{n}+x_{n}g_{n})$ induces a permutation of $R^{n}$, that is $\pi_{n}(\vec{f})$ is a permutation of $R^{n}$.
Proof.
Set $\vec{F}=\pi_{n}(\vec{f})$. First, we show that $\vec{F}$ is one-to-one. For this, let $a_{i},b_{i}\in R$ for $i=1,\dots,n$ such that
$\vec{F}(a_{1},\ldots,a_{n})=\vec{f}(a_{1},\ldots,a_{n})=\vec{f}(b_{1},\ldots,b_{n})=\vec{F}(b_{1},\ldots,b_{n}).$
Then, we have the following equations:
$\begin{matrix}f_{1}(a_{1})=f_{1}(b_{1})\\ f_{2}(a_{1})+a_{2}g_{1}(a_{1})=f_{2}(b_{1})+b_{2}g_{1}(b_{1})\\ \vdots\\ f_{n}(a_{1},\ldots,a_{n-1})+a_{n}g_{n}(a_{1},\ldots,a_{n-1})=f_{n}(b_{1},\ldots,b_{n-1})+b_{n}g_{n}(b_{1},\ldots,b_{n-1})\end{matrix}$
Thus, from the first equation, since $f_{1}$ is a permutation polynomial on $R$, we have $a_{1}=b_{1}$. Then, the second equation becomes $a_{2}g_{2}(a_{1})=b_{2}g_{2}(a_{1})$. But, then $a_{2}=b_{2}$ as $g_{2}(a_{1})$ is a unit since $g_{2}$ is a unit-valued polynomial on $R$. Continuing in this manner, we find that $a_{i}=b_{i}$, $11\leq i\leq n-$ 1 and the $i+1$th equation is reduced to $a_{i+1}g_{i+1}(a_{1},\ldots,a_{i})=b_{i+1}g_{i+1}(a_{1},\ldots,a_{i})$. Hence, $a_{i+1}=b_{i+1}$ since $g_{i+1}$ is a unit-valued polynomial. Therefore, $a_{i}=b_{i}$ for $i=1,\dots,n$. To show $\vec{F}$ is surjective, let $(c_{1},\ldots,c_{n})\in R^{n}$ be arbitrary. Then, we solve the equation
$\vec{f}(x_{1},\ldots,x_{n})=(c_{1},\ldots,c_{n}).$
Therefore, we have the following system of $n$ equations:
$\begin{matrix}f_{1}(x_{1})=c_{1}\\ f_{2}(x_{1})+x_{2}g_{2}(x_{1})=c_{2}\\ \vdots\\ f_{n}(x_{1},\ldots,x_{n-1})+x_{n}g_{n}(x_{1},\ldots,x_{n-1})=c_{n}.\end{matrix}$
Since $f_{1}$ is a permutation on $R$, $x_{1}=a_{1}=F_{1}^{-1}(c_{1})$ is the solution of the 1st equation, where $F_{1}^{-1}$ stands for the inverse permutation of the permutation obtained by $f_{1}$ on $R$. So, the second equation becomes $f_{2}(a_{1})+x_{2}g_{2}(a_{1})=c_{2}$. Since $g_{2}$ is a unit-valued on $R$, $g_{2}(a_{1})$ is invertible in $R$. Thus, $x_{2}=a_{2}=(c_{2}-f_{2}(a_{1})){g_{2}(a_{1})}^{-1}$. Solving these equations and the fact that $g_{i}$ is a unit valued on $R$ yield
$x_{i}=a_{i}=(c_{i-1}-f_{i}(a_{1},\ldots,a_{i-1})){g_{i}(a_{1},\ldots,a_{i-1})}^{-1}\text{, }i=2,\ldots,n.\qed$
Theorem 3.3.
Let $R$ be a commutative ring with unity and $n>1$ an integer. Let $\mathcal{MT}_{n}$ be the set of all vector-polynomials of the form:
$\vec{f}(x_{1},\ldots,x_{n})=(f_{1}(x_{1}),f_{2}(x_{1})+x_{2}g_{2}(x_{1}),\ldots,f_{n}(x_{1},\ldots,x_{n-1})+x_{n}g_{n}(x_{1},\ldots,x_{n-1})),$ (3)
where $f_{1}$ is a permutation polynomial on $R$, $f_{i}\in R_{[{i-1}]}$ and $g_{i}\in UV(R_{[{i-1}]]})$ for $i=2,\ldots,n$. Then $\mathcal{MT}_{n}$ is a monoid of vector-permutation polynomials with respect to composition. Moreover, if $R$ is finite, then $\pi_{n}(\mathcal{MT}_{n})$ is a group of permutations of $R^{n}$.
Proof.
By Lemma 3.2, $\mathcal{MT}_{n}$ consists of vector-permutation polynomials. To show $\mathcal{MT}_{n}$ is a monoid with respect to the composition ``"``"`$\circ$"` ` ∘ " we only need to show that $\mathcal{MT}_{n}$ is closed with respect to ``"``"`$\circ$"` ` ∘ " since the identity $(x_{1},x_{2},\ldots,x_{n})\in\mathcal{MT}_{n}$ and composition is always associative (see Definition 2.6). But, this is the case by Lemma 3.1.
Moreover, if $S_{R^{n}}$ is the symmetric group on $R^{n}$, then the first part of the proof shows that $\pi_{n}(\mathcal{MT}_{n})\subseteq S_{R^{n}}$. Since we consider the case $R$ is finite, $\pi_{n}(\mathcal{MT}_{n})$ is a finite subset of the finite group $S_{R^{n}}$. Thus, we need only show that $\pi_{n}(\mathcal{MT}_{n})$ is closed with respect to composition. Now, let $\vec{F},\vec{G}\in\pi_{n}(\mathcal{MT}_{n})$. Then there exist $\vec{f},\vec{g}\in\mathcal{MT}_{n}$ such that $\vec{F}=\pi_{n}(\vec{f})$ and $\vec{G}=\pi_{n}(\vec{g})$. Therefore,
$\vec{F}\circ\vec{G}=\pi_{n}(\vec{f})\circ\pi_{n}(\vec{g})=\pi_{n}(\vec{f}\circ\vec{g})\in\pi_{n}(\mathcal{MT}_{n}).\qed$
As a direct result, we have the following corollary.
Corollary 3.4.
Let $R$ be a finite commutative ring. Let $k>1$ and $f,g\in R_{[k]}$. Suppose that $f$ is unit-valued. Then $g(x_{1},\ldots,x_{k})+x_{k+1}f(x_{1},\ldots,x_{k})$ is a permutation polynomial of $i$ variables for every $i\geq k+1$.
Remark 3.5.
The finiteness condition on $R$ in the last part of Theorem 3.3 is sufficient for $\pi_{n}(\mathcal{MT})_{n}$ to be a group. However, if we consider the case in which $R$ is an infinite field, then $\pi_{n}(\mathcal{MT}_{n})\cong\mathcal{MT}_{n}$ because if $f\in R_{[k]}$ ($k\geq 1$) maps $R^{k}$ onto zero then $f$ is necessarily the constant zero polynomial. Therefore, we notice the following examples:
  1. (1)
    Let $R$ be a real closed field (for example the field of real numbers and the field of real algebraic numbers see [27]). Then $x^{2k+1}$ is a permutation polynomial on $R$ for every $k\geq 1$ by [4, Theorem 9]. So, it is evident that $(x_{1}^{2k+1},x_{2},x_{3},\ldots,x_{n})$ is a non-invertible element in $\mathcal{MT}_{n}$ for $k\geq 1$. Hence, $\pi_{n}(\mathcal{MT}_{n})$ is not a group since $\pi_{n}((x_{1}^{2k+1},x_{2},x_{3},\ldots,x_{n}))$ is a non-invertible element in $\pi_{n}(\mathcal{MT}_{n})$ for $k\geq 1$.
  2. (2)
    Let $R$ be an algebraically closed field of characteristic zero. Then $\pi_{n}(\mathcal{MT}_{n})$ is an infinite group, because in this case, $\mathcal{MT}_{n}$ is a group (we see this later in Proposition 4.12).
  3. (3)
    Let $R$ be an algebraically closed field of characteristic $p$. Then $ax^{p^{t}}$ is a permutation polynomial on $R$ for every $a\neq 0$ and $t\geq 1$ by [5, Theorem 7]. Hence, $(ax_{1}^{p^{t}},x_{2},x_{3},\ldots,x_{n})$ is a non-invertible element in $\mathcal{MT}_{n}$ for every $a\neq 0$ and $t\geq 1$. Therefore, $\pi_{n}(\mathcal{MT}_{n})$ is not a group since $\pi_{n}((ax_{1}^{p^{t}},x_{2},x_{3},\ldots,x_{n}))$ is a non-invertible element in $\pi_{n}(\mathcal{MT}_{n})$ provided that $a\neq 0$ and $t\geq 1$.
Definition 3.6.
Let $n>1$. We call the monoid $\mathcal{MT}_{n}$ constructed in Theorem 3.3 the monoid of triangular vector-permutation polynomials or more briefly the triangular monoid.
Notation 3.7.
We set $\mathcal{MT}_{1}$ to be the monoid $\mathcal{MP}(R)$ of permutation polynomials on $R$.
Our next aim is to describe the group of units of the monoid of vector-permutation polynomials constructed in Theorem 3.3. Before doing so, we need to fix some notation.
Notation 3.8.
  1. (1)
    Let $(1:f)$ denote $(f(x_{1}),x_{2},\ldots,x_{n})$, where $f\in R[x_{1}]$ is a permutation polynomial on $R$.
  2. (2)
    For $i>1$, let $(i:u;f)$ denote $(x_{1},\ldots,x_{i-1},f+x_{i}u,x_{i+1},\ldots,x_{n})$, where $f,u\in R_{[i-1]}$ such that $u\in UV(R_{[i-1]})$ (i.e. $u$ is a unit-valued polynomial).
  3. (3)
    When there is no confusion we write $\vec{f}\vec{g}$ instead of $\vec{f}\circ\vec{g}$.
Remark 3.9.
  1. (1)
    Let $\vec{f}\in\mathcal{MT}_{n}$. By definition, $\vec{f}$ has the form  (3) in Theorem 3.3. Then in view of Notation 3.8, we have by Fact 2.7,
    $\vec{f}=(1:f_{1})\circ(2:u_{2};f_{2})\circ\cdots\circ(n:u_{n};f_{n}),\text{or simply we write}$
    $\vec{f}=(1:f_{1})\prod_{i=2}^{n}(i:u_{i};f_{i}),$ (4)
    where $f_{1}\in R[x_{1}]$ is a permutation polynomial and $f_{i},u_{i}\in R_{[i-1]}$ such that $u_{i}$ is a unit-valued polynomial for $i=2,\ldots,n$. Clearly, that the expression  (4) is uniquely determined by the expression (3)
  2. (2)
    Again in view of Notation 3.8, we see that for $i>1$, $(1:x_{1})=(i:1;0)=(x_{1},\ldots,x_{n})$ is the identity element of $\mathcal{MT}_{n}$.
  3. (3)
    Let $\vec{g}=(1:g_{1})\prod_{i=2}^{n}(i:w_{i};g_{i})$. To obtain an expression of $\vec{g}\vec{f}$ similar to expression (4), we adopt Equations (1) and  (2) to get
    $\vec{g}\vec{f}=(1:g_{1}\circ f_{1})\prod_{i=2}^{n}(i:u_{i}w_{i}(\vec{v_{i}});g_{i}(\vec{v_{i}})+w_{i}(\vec{v_{i}})f_{i}),\text{ where }$ (5)
    $\vec{v_{i}}(x_{1},\ldots,x_{i-1})$ $=(f_{1}(x_{1}),f_{2}(x_{1})+x_{2}u_{2}(x_{1}),\ldots,f_{i-1}(x_{1},\ldots,x_{i-1})+x_{{i-1}}u_{i-1}(x_{1},\ldots,x_{i-2}));$
    $\quad i=2,\ldots,n.$
    We notice here that $w_{i}(\vec{v_{i}})$ is a unit-valued polynomial since $w_{i}$ is a unit-valued polynomial. So, expression (5) can be written in the expression  (4) by choosing $h_{1}=g_{1}\circ f_{1}$,
The following Lemmas are straightforward.
Lemma 3.10.
Let $f,g\in R[x_{1}]$ be permutation polynomials. Then
  1. (1)
    $(1:f)(1:g)=(1:f\circ g)$;
  2. (2)
    $\vec{f}=(1:f)$ is a unit (invertible) in $\mathcal{MT}_{n}$ if and only if $f$ is invertible (with respect to the composition of polynomials) if and only if $f$ is an $R$-automorphism of $R[x_{1}]$. In this case $\vec{f}^{-1}=(1:f^{-1})$, where $f^{-1}$ is the inverse of $f$ with respect to composition.
Lemma 3.11.
Let $i>1$ and $f,g,u,v\in R_{[i-1]}$ such that $u$ and $v$ are unit-valued polynomials. Then
  1. (1)
    $(i:u;f)(i:v;g)=(i:uv;f+ug)$;
  2. (2)
    $\vec{f}=(i\!:u;f)$ is a unit in $\mathcal{MT}_{n}$ if and only if $u$ is a unit of the polynomial ring $R_{[i-1]}$ (with respect to multiplication). In this case, $\vec{f}^{-1}=(i:u^{-1};-u^{-1}f)$, where $u^{-1}$ is the inverse of $u$ in the ring $R_{[i-1]}$ (with respect to multiplication).
By means of Notation 3.8, Remark 3.9, Definition 2.8 and Lemma 2.13, we get the following useful lemma.
Lemma 3.12.
Let $R$ be a finite commutative ring. Let $f_{1},g_{1}\in R[x_{1}]$ be permutation polynomials on $R$, and $f_{i},g_{i},v_{i},u_{i}\in R_{[i-1]}$ such that $u_{i}$ and $v_{i}$ are unit-valued polynomials on $R$ for $i=2,\ldots,n$. The following statements are equivalent
  1. (1)
    $(1:f_{1})\prod_{i=2}^{n}(i:u_{i};f_{i})\operatorname{{}\triangleq{}}(1:g_{1})\prod_{i=2}^{n}(i:v_{i};g_{i})$ on $R^{n}$;
  2. (2)
    $(1:f_{1})\operatorname{{}\triangleq{}}(1:g_{1})$ and $(i:u_{i};f_{i})\operatorname{{}\triangleq{}}(i:v_{i};g_{i})$ on $R^{n}$ for $i=2,\dots,n$;
  3. (3)
    $f_{1}\operatorname{{}\triangleq{}}g_{1}$, $f_{i}\operatorname{{}\triangleq{}}g_{i}$ and $u_{i}\operatorname{{}\triangleq{}}v_{i}$ on $R$ for $i=2,\dots,n$.
Theorem 3.13.
Let $R$ be a commutative ring. Then the group of units of the monoid of triangular vector-permutation polynomials $\mathcal{MT}_{n}$ consists of all elements of the form
$\vec{f}=(1:f_{1})\prod_{i=2}^{n}(i:u_{i};f_{i}),$
where $f_{1}\in R[x_{1}]$ is an $R$-automorphism, $f_{i}\in R_{[i-1]}$ and $u_{i}\in R_{[i-1]}^{\times}$ for $i=2,\ldots,n$.
Proof.
Let $\vec{f}\in\mathcal{MT}_{n}$. Then, by Remark 3.9, $\vec{f}=(1:f_{1})\prod_{i=2}^{n}(i:u_{i};f_{i})$ where $f_{1}\in\mathcal{MP}(R)$, $f_{i}\in R_{[i-1]}$ and $u_{i}\in UV(R_{[i-1]})$. First, suppose that $\vec{f}$ is a unit in $\mathcal{MT}_{n}$. Then there exists $\vec{g}=(1:g_{1})\prod_{i=2}^{n}(i:w_{i};g_{i})$, where $g_{1}\in\mathcal{MP}(R)$, $g_{i}\in R_{[i-1]}$ and $w_{i}\in UV(R_{[i-1]})$ with $\vec{g}\vec{f}=(x_{1},\ldots,x_{n})$. By Remark 3.9 (Equation (5)),
$\vec{g}\vec{f}=(1:g_{1}\circ f_{1})\prod_{i=2}^{n}(i:u_{i}w_{i}(\vec{v_{i}});g_{i}(\vec{v_{i}})+w_{i}(\vec{v_{i}})f_{i}),\text{ where }\vec{v_{i}}\in R_{[i-1]}^{i-1}\text{ for }i=2,\ldots,n.$
Again by Remark 3.9, we have the following equalities
$\begin{matrix}g_{1}\circ f_{1}=x_{1}\\ g_{i}(\vec{v_{i}})+w_{i}(\vec{v_{i}})f_{i}x_{i}=x_{i}\text{ for }i=2,\ldots,n.\end{matrix}$
The first equality implies that $f_{1}$ is an $R$-automorphism. From the second equality, we infer that $w_{i}(\vec{v_{i}})f_{i}=1$ since $g_{i}(\vec{v_{i}}),w_{i}(\vec{v_{i}})$ and $f_{i}$ are independent from the variable $x_{i}$ for $i=2,\ldots,n$. Thus, $f_{i}\in R_{[i-1]}^{\times}$. Conversely, assume that $f_{1}\in R[x_{1}]$ is an $R$-automorphism, $f_{i}\in R_{[i-1]}$ and $u_{i}\in R_{[i-1]}^{\times}$ for $i=2,\ldots,n$. Then by construction and Lemma 3.10 and Lemma 3.11, $\vec{f}$ is invertible being product (composition) of units in $\mathcal{MT}_{n}$. ∎
From Theorem 3.13, Lemma 3.10 and Lemma 3.11 we have the following result.
Corollary 3.14.
Let $\vec{f}\in\mathcal{MT}_{n}$. Suppose that $\vec{f}=(1:f_{1})\prod_{i=2}^{n}(i:u_{i};f_{i})$ where $f_{0}\in\mathcal{MP}(R)$, $f_{i},u_{i}\in R_{[i-1]}$ such that $u_{i}\in UV(R_{[i-1]})$. Then $\vec{f}$ is invertible if and only if $(1:f_{1})$ and $(i:u_{i};f_{i})$ are invertible for $i=2,\ldots,n$. Further, the inverse is given by the relation
$\vec{f}^{-1}=(\prod_{i=n}^{2}(i:u_{i}^{-1};-u_{i}^{-1}f_{i}))(1:f_{1}^{-1}),$
where $f_{1}^{-1}$ is the inverse of $f_{1}$ with respect to composition and for $i=2,\ldots,n$, $u_{i}^{-1}$ is the inverse of $u_{i}$ with respect to multiplication.
Definition 3.15.
We denote by $TR_{n}$ the group of units of the monoid $\mathcal{MT}_{n}$. We call $TR_{n}$ the group of triangular vector-permutation polynomials or shortly the triangular group.
If $R$ is finite, then $\pi_{n}({\mathcal{MT}_{n}})$ is a group of permutations of $R^{n}$ by Theorem 3.13. Furthermore, in this case, $\pi_{n}({TR_{n}})$ is also a group of permutations of $R^{n}$ (being a closed subset of the finite group $\pi_{n}({\mathcal{MT}_{n}})$). An interesting question arises from what we already mentioned: Is it true that $\pi(\mathcal{MT}_{n})\neq\pi_{n}(TR_{n})$? Fortunately, we can generally answer this question affirmatively except in the case in which $R$ is a finite product of the field of two elements $\mathbb{F}_{2}$. Without loss of generality, we consider only the case that $R$ is a finite local ring since the general case follows by the Chinese Reminder Theorem.
Before doing so, we recall the following celebrated fact which is a special case of a more general result due to Nöbauer [19, Theorem 2.3]. For a proof of this fact, we refer the interested reader to a paper by Nechaev [17, Theorem 3].
Hereafter, for a polynomial $f$, let $f^{\prime}$ be its formal derivative.
Lemma 3.16.
[19, Theorem. 2.3] Let $R$ be a finite local ring, that is not a field, $M$ its maximal ideal, and $f\in R[x]$.
Then $f$ is a permutation polynomial on $R$ if and only if the following conditions hold:
  1. (1)
    $f$ is a permutation polynomial on $R/M$;
  2. (2)
    for all $a\in R$, $f^{\prime}(a)\neq 0\mod{M}$.
Lemma 3.17.
Let $R$ be a finite local ring which is not a field with residue field $\mathbb{F}_{q}$. Then there is no invertible unit-valued polynomial that represents the unit-valued polynomial function induced by $g(x)=(x^{q}-x)+1$, i.e. $\pi_{1}(g)\in\pi_{1}(UV(R[x])\setminus{R[x]}^{\times})$.
Proof.
Assume to the contrary that there is an invertible $f\in R[x]$ such that $f$ and $g$ induce the same function on $R$. Then by Remark 2.4, since $f(0)=g(0)=1$, $f(x)=1+m_{1}x+\cdots+m_{n}x^{n}$ with $m_{1},\ldots,m_{n}\in M$. Also, $xf(x)$ and $xg(x)$ induce the same function on $R$. However, by Lemma 3.16, $xf(x)$ is a permutation polynomial while $xg(x)$ is not (since the derivative of $xg(x)$ admits a root on $\mathbb{F}_{q}$). This contradiction implies that such a polynomial $f$ does not exist. ∎
Theorem 3.18.
Let $R$ be a finite local ring. Then $\pi_{n}(TR_{n})=\pi_{n}(\mathcal{MT}_{n})$ if and only if $R=\mathbb{F}_{2}$.
Proof.
If $R=\mathbb{F}_{2}$, one easily sees that $1$ is the only unit of $R$ and hence every unit-valued polynomial $u(x_{1},$$,x_{k})$ in $k$ variables (for any $k\geq 1$) is equivalent to the (invertible) constant polynomial $1$. On the other hand, there are only two permutations of $R$ that are induced by the (invertible) permutation polynomials $x$ and $x+1$.
Now, if $\vec{F}\in\pi_{n}(\mathcal{MT}_{n})$, then $\vec{F}=\pi_{n}(\vec{f})$ for some $\vec{f}\in\pi(\mathcal{MT}_{n})$, where
$\vec{f}=(1:f_{1})\prod_{i=2}^{n}(i:u_{i};f_{i}),$
with $f_{1}\in\mathcal{MP}(R)$ and $f_{i},u_{i}\in R_{[i-1]}$ such that $u_{i}\in UV(R_{[i-1]})$ for $i=2,\ldots,n$. But, the above discussion tells us that $u_{i}\operatorname{{}\triangleq{}}1$ on $R$ (for $i=2,\ldots,n$) and either $f_{1}\operatorname{{}\triangleq{}}x_{1}$ or $x_{1}+1$ on $R$. Without loss of generality say $f_{1}\operatorname{{}\triangleq{}}x_{1}$. So, if we consider $\vec{g}=(1:x_{1})\prod_{i=2}^{n}(i:1;f_{i})$, then, by Theorem 3.13 and Definition 2.8, $\vec{g}\in TR_{n}$ and $\pi_{n}(\vec{g})=\pi_{n}(\vec{f})=\vec{F}$. Thus, $\vec{F}\in\pi_{n}(TR_{n})$.
For the other direction, we construct a vector-polynomial permutation $\vec{F}\in\pi_{n}(\mathcal{MT}_{n})\setminus\pi_{n}(TR_{n})$. However, we distinguish two cases on $R$. First, assume that $R$ is a local ring with maximal ideal $M\neq\{0\}$, and consider the unit-valued polynomial $g(x_{1})=(x_{1}^{q}-x_{1})+1$. Then $\vec{g}=(2:g;0)\in\mathcal{MT}_{n}\setminus TR_{n}$ and hence $\pi(\vec{g})\in\pi(\mathcal{MT}_{n})$. We claim that $\pi(\vec{g})\notin\pi(TR_{n})$. Otherwise, suppose that there is a vector-permutation polynomial $\vec{f}\in TR_{n}$, such that $\pi_{n}(\vec{f})=\pi_{n}(\vec{g})$. Then, writing $\vec{f}=(1:f_{1})\prod_{i=2}^{n}(i:u_{i};f_{i})$ implies that $u_{2}$ is a unit of $R[x_{1}]$ by Theorem 3.13. Since $\pi(\vec{g})=\pi_{n}(\vec{f})$, we have by Lemma 3.12, $u_{2}\operatorname{{}\triangleq{}}g$ on $R$. But, such a unit polynomial $u_{2}$ does not exist by Lemma 3.17. Hence, the claim.
Now, consider the case $R=\mathbb{F}_{q}$ with $q\neq 2$. One can easily see that the unit polynomials are only the non zero-constants. Then, since $q\neq 2$, there is an element $a\in\mathbb{F}_{q}$ such that $a\neq 0,1$ and we can define a non constant unit-valued function $F\colon\mathbb{F}_{q}\longrightarrow\mathbb{F}_{q}$, $F(b)=\begin{cases}a&\textnormal{if }b=a\\ 1&\textnormal{if }b\neq a\end{cases}.$ Also, by Lagrange interpolation, we can find a polynomial $g$ that represents $F$. Then an argument similar to that of the previous case shows that $\pi_{n}(2:g;0)\in\pi_{n}(\mathcal{MT}_{n})\setminus\pi_{n}(TR_{n})$. ∎
Remark 3.19.
Let $k>n$. Then the map
$e_{n,k}\colon R_{[n]}^{n}\longrightarrow R_{[k]}^{k},\vec{f}=(f_{1},\ldots,f_{n})\mapsto(f_{1},\ldots,f_{n},x_{n+1}\ldots,x_{k})$
is well known to be an embedding and by its restrictions on the monoid $\mathcal{MT}_{n}$ and its group of units $TR_{n}$ respectively, we have the following embeddings
  1. (1)
    $\mathcal{MT}_{n}\hookrightarrow\mathcal{MT}_{k}$; and
  2. (2)
    $TR_{n}\hookrightarrow TR_{k}$.
Also, there is an embedding $\bar{e}_{n,k}\colon\pi_{n}(\mathcal{MT}_{n})\longrightarrow\pi_{k}(\mathcal{MT}_{k}),\pi_{n}(\vec{f})\mapsto\pi_{k}\circ e_{n,k}(\vec{f})$.
4. Some results on the structure of $\mathcal{MT}_{n}$
In this section, we investigate the structures of the monoid $\mathcal{MT}_{n}$, its group of units $TR_{n}$, and when $R$ is finite, its induced group $\pi_{n}(\mathcal{MT}_{n})$ of permutations on $R^{n}$. In particular, we see that the structures of $TR_{n}$ and $\pi_{n}(\mathcal{MT}_{n})$ depend mainly on that of $\mathcal{MT}_{n}$.
From now on, when $H$ is a monoid let $H^{\times}$ denote its group of units.
4.1. The structure of the monoid $\mathcal{MT}_{n}$
In this subsection, we see that the triangular monoid $\mathcal{MT}_{n}$ can be decomposed into an iterated semi-direct product of $n$ monoids.
The concept of the semi-direct product of monoids plays a vital role in this part of the paper. In the following, we recall its definition. Let $A$ and $B$ be monoids and $End(B)$ be the monoid of endomorphisms of $B$ with respect to composition. If $\phi\colon A\longrightarrow End(B)$, $a\mapsto\phi_{a}$, is a homomorphism then the semi-direct product $B\rtimes_{\phi}A$ (or simply $B\rtimes A$) is the monoid with elements $\{(a,b)\colon a\in A,b\in B\}$ and operation $(a,b)(c,d)=(ac,b\phi_{a}(d))$ (when $A$ acts on $B$ from left) or $(a,b)(c,d)=(ac,\phi_{c}(b)d)$ (when $A$ acts on $B$ from right) (see for example [18]). We note to the reader that if we chose $A$ and $B$ to be two groups, and replace $End(B)$ by the group $Aut(B)$ of automorphisms (i.e. $\phi\colon A\longrightarrow Aut(B)$), then we just have the semi-direct product of the group $B$ by $A$ with the same operation as mentioned before. We note here that $(a,b)^{-1}=(a^{-1},\phi_{a^{-1}}(b^{-1}))$ and $\phi_{a^{-1}}=\phi_{a}^{-1}$.
Lemma 4.1.
Fix $2\leq k\leq n$. Let $\mathcal{ML}^{n}_{k}$ denote the set of vector-polynomials of the form
$(k:u;f)$, where $f,u\in R_{[k-1]}$ such that $u\in UV(R_{[k-1]})$, is a submonoid of the monoid $\mathcal{MT}_{n}$ which is isomorphic to the semi-direct product of the additive group of $R_{[k-1]}$ by the monoid $(UV(R_{[k-1]}),``\cdot")$ of unit-valued polynomials. That is, $\mathcal{ML}^{n}_{k}\cong R_{[k-1]}\rtimes UV(R_{[k-1]})$.
Proof.
It is clear that the identity $(k:1;0)\in\mathcal{ML}^{n}_{k}$. Also, by Lemma 3.11, $\mathcal{ML}^{n}_{k}$ is closed with respect to composition, whence $\mathcal{ML}^{n}_{k}$ is a submonoid of $\mathcal{MT}_{n}$. Now, given a unit-valued polynomial $u\in UV(R_{[k-1]})$, one can define an endomorphism $\phi_{u}\colon R_{[k-1]}\longrightarrow R_{[k-1]}$ by $\phi_{u}(f)=uf$ for every $f\in R_{[k-1]}$. Consider, the map $\phi\colon UV(R_{[k-1]})\longrightarrow(End(R_{[k-1]}),``\circ")$, ($u\mapsto\phi_{u}$). We claim that $\phi$ is a homomorphism. To show the claim, let $u,v\in UV(R_{[k-1]})$ and consider $f\in R_{[k-1]}$. We have,
$\phi_{u}\circ\phi_{v}(f)=\phi_{u}(\phi_{v}(f))=\phi_{u}(vf)=uvf=\phi_{uv}(f).$
Thus $\phi_{u}\circ\phi_{v}=\phi_{uv}.$ This proves the claim. Hence, the semi-direct product $R_{[k-1]}\rtimes UV(R_{[k-1]})$ is defined with the following operation
$(u,f)(v,g)=(uv,f+ug),$
where $u,v\in UV(R_{[k-1]})$ and $f,g\in R_{[k-1]}$. The rest is to show that the map
$\psi\colon R_{[k-1]}\rtimes UV(R_{[k-1]})\longrightarrow\mathcal{ML}^{n}_{k},\text{ }(u,f)\mapsto(k:u;f)$
is an isomorphism. We leave this to the reader. ∎
Next, we show that the monoid $\mathcal{MT}_{n}$ can be viewed as the semi-direct product of the $\mathcal{ML}^{n}_{n}$ (defined in Lemma 4.1) by the monoid $\mathcal{MT}_{n-1}$. We begin with proving the following technical Lemma.
Lemma 4.2.
Let $n>1$. Then $\mathcal{ML}^{n}_{n}\rtimes\mathcal{MT}_{n-1}$ is equipped with the operation
$(\vec{h},((n:u;f))(\vec{l},((n:v;g))=(\vec{h}\circ\vec{l},(n:u(\vec{l})v;f(\vec{l})+u(\vec{l})g)),$ (6)
where $\vec{h},\vec{l}\in\mathcal{MT}_{n-1}$ and $(n:u;f),(n:v;g)\in\mathcal{ML}^{n}_{n}$
Proof.
Let $\vec{h}\in\mathcal{MT}_{n-1}$. For simplicity write $\vec{h}=(h_{1},\ldots,h_{n-1})$, where $h_{i}\in R_{[i]}$ for $i=1,\ldots,n-1$ (keeping in mind that $\vec{h}$ satisfies the conditions of expression (3) in Theorem 3.3). Then define a map $\phi_{\vec{h}}\colon\mathcal{ML}^{n}_{n}\longrightarrow\mathcal{ML}^{n}_{n}$ by $((n:u;f))\phi_{\vec{h}}=(n:u(\vec{h});f(\vec{h}))$ for $(n:u;f)\in\mathcal{ML}^{n}_{n}$. Since $u$ is a unit-valued polynomial (in $n-1$ variables), we have that $u(\vec{h})$ is a unit-valued polynomial (in $n-1$ variables). This shows $((n:u;f))\phi_{\vec{h}}\in\mathcal{ML}^{n}_{n}$. Now, let $(n:v;g)\in\mathcal{ML}^{n}_{n}$ and consider
[…]
$=(n:(uv)(\vec{h});f(\vec{h})+(ug)(\vec{h}))=(n:u(\vec{h})v(\vec{h});f(\vec{h})+u(\vec{h})g(\vec{h}))$
[…]
Thus, $\phi_{\vec{h}}\in End(\mathcal{ML}^{n}_{n})$. Then, the map $\phi\colon\mathcal{MT}_{n-1}\longrightarrow End(\mathcal{ML}^{n}_{n})$, ${\vec{h}}\mapsto\phi_{\vec{h}}$ is a homomorphism. Indeed, if $\vec{l}=(l_{1},\ldots,l_{n-1})\in\mathcal{MT}_{n-1}$, we have
$((n:u;f))\phi_{\vec{h}\circ\vec{l}}$ $=(n:u(\vec{h}\circ\vec{l});f(\vec{h}\circ\vec{l}))=(n:u(h_{1}(\vec{l}),\ldots,h_{n-1}(\vec{l}));f(h_{1}(\vec{l}),\ldots,h_{n-1}(\vec{l})))$
$=(n:u(h_{1},\ldots,h_{n-1});f(h_{1},\ldots,h_{n-1}))\phi_{\vec{l}}=((n:u;f))\phi_{\vec{h}})\phi_{\vec{l}}$
$=((n:u;f))\phi_{\vec{h}}\circ\phi_{\vec{l}}.$
Thus, the semi-direct product $\mathcal{ML}^{n}_{n}\rtimes\mathcal{MT}_{n-1}$ is defined with the following operation
$(\vec{h},((n:u;f))(\vec{l},((n:v;g))$ $=(\vec{h}\circ\vec{l},((n:u;f))\phi_{\vec{l}}(n:v;g))$
$=(\vec{h}\circ\vec{l},(n:u(\vec{l})v;f(\vec{l})+u(\vec{l})g))$
or more explicitly
$(\vec{h},(n:u;f))(\vec{l},(n:v;g))$ $=((h_{1}\circ l_{1},h_{2}(l_{1},l_{2}),\ldots,h_{n-1}(l_{1},\ldots,l_{n-1})),(n:u(\vec{l})v;f(\vec{l})+u(\vec{l})g)).$ (7)
Proposition 4.3.
Let $n>1$ and let $\mathcal{ML}^{n}_{n}$ denote the monoid of vector-polynomials of the form
$(n:u;f)$, where $f,u\in R_{[k-1]}$ and $u\in UV(R_{[k-1]})$. Then $\mathcal{MT}_{n}\cong\mathcal{ML}^{n}_{n}\rtimes\mathcal{MT}_{n-1}$.
Proof.
In view of Lemma 4.2, the semi-direct product $\mathcal{ML}^{n}_{n}\rtimes\mathcal{MT}_{n-1}$ is defined by the operation given in Equation (6) (more explicit in Equation (7)). Define a map $\psi\colon\mathcal{ML}^{n}_{n}\rtimes\mathcal{MT}_{n-1}\longrightarrow\mathcal{MT}_{n}$ as the following: if $(\vec{h},(n:u;f))\in\mathcal{ML}^{n}_{n}\rtimes\mathcal{MT}_{n-1}$, where $u\in UV(R_{[n-1]})$ and $\vec{h}=(h_{1},\ldots,h_{n-1})$; $h_{i}\in R_{[i]}$ for $i=1,\ldots,n-1$, then set
$\psi(\vec{h},(n:u;f))=(h_{1},\ldots,h_{n-1},f+ux_{n}).$
Note that, by the definition of $\mathcal{MT}_{n-1}$, $h_{1}$ is a permutation polynomial on $R$ and $h_{i}=f_{i}+u_{i}x_{i}$ where $f_{i},\in R_{[i-1]}$ and $u_{i}\in UV(R_{[i-1]})$ for $i=2,\ldots,n-1$. This means that $\psi(\vec{h},(n:u;f))\in\mathcal{MT}_{n}$. Now, we show that $\psi$ is a homomorphism. Let $(\vec{h},(n:u;f))$ and $(\vec{l},(n:v;g))\in\mathcal{ML}^{n}_{n}\rtimes\mathcal{MT}_{n-1}$. Then we have
$)\displaystyle\psi((\vec{h},(n:u;f)$$)(\vec{l},(n:v;g)))$
[…]
$=(h_{1}\circ l_{1},h_{2}(l_{1},l_{2}),\ldots,h_{n-1}(l_{1},\ldots,l_{n-1}),f(\vec{l})+u(\vec{l})g+u\big{(}\small{\vec{l}}\big{)}vx_{n})$
$=(h_{1}\circ l_{1},h_{2}(l_{1},l_{2}),\ldots,h_{n-1}(l_{1},\ldots,l_{n-1}),f(l_{1},\ldots,l_{n-1})+u(l_{1},\ldots,l_{n-1})g$
$\quad+u(l_{1},\ldots,l_{n-1})vx_{n})$
$=(h_{1},\ldots,h_{n-1},f+ux_{n})(l_{1},\ldots,l_{n-1},g+vx_{n})$
$=\psi((\vec{h},(n:u;f)))\psi((\vec{l},(n:v;g))).$
The hardest part of the proof has been finished, and the remaining details are straightforward and left to the reader. ∎
Convention 4.4.
From now on, by $A_{1}\rtimes A_{2}\rtimes A_{3}$, we mean $A_{1}\rtimes(A_{2}\rtimes A_{3})$, i.e. we start computing from the left.
We are now in position to decompose the triangular monoid $\mathcal{MT}_{n}$ into an iterated semi-direct product of $\mathcal{ML}^{n}_{n},\ldots,\mathcal{ML}^{2}_{2}$, and $\mathcal{MT}_{1}=\mathcal{MP}(R)$ (see Notation 3.7). In fact, we deduce more.
Theorem 4.5.
Let $n>1$. Then
  1. (1)
    $\mathcal{MT}_{n}\cong\mathcal{ML}^{n}_{n}\rtimes\cdots\rtimes\mathcal{ML}^{2}_{2}\rtimes\mathcal{MP}(R)$;
  2. (2)
    $\mathcal{MT}_{n}\cong(R_{[n-1]}]\rtimes UV(R_{[n-1]}))\rtimes\cdots\rtimes(R{[x_{1}]}\rtimes UV(R{[x_{1}]}))\rtimes\mathcal{MP}(R)$.
Proof.
By Proposition 4.3, $\mathcal{MT}_{n}\cong\mathcal{ML}^{n}_{n}\rtimes\mathcal{MT}_{n-1}$. Then iterated applications of this relation yield the first expression. Then, we use Lemma 4.1, to have $\mathcal{ML}^{i}_{i}\cong R_{[i-1]}\rtimes UV(R_{[i-1]})$ for $i=2,\ldots,n$. Combining this with the first expression gives the second expression. ∎
4.2. The structure of the triangular group $TR_{n}$
Using the fact that monoid homomorphisms preserve the identity, we will see that the group units of a semi-direct product of monoids is just the semi-direct product of their groups of units. Then, we apply this result to the decomposition of $\mathcal{MT}_{n}$ given in Theorem 4.5 to obtain a decomposition of the group $TR_{n}$ into an iterated semi-direct product of groups.
Lemma 4.6.
Let $A$ and $B$ be monoids. Consider the homomorphism $\phi\colon A\longrightarrow End(B)$ ($a\mapsto\phi_{a}$, $\phi_{a}\colon B\longrightarrow B$ is a homomorphism), and let $\psi_{a}$ be the restriction of $\phi_{a}$ onto $B^{\times}$ then
  1. (1)
    $\phi_{a}\in Aut(B)$ for every $a\in A^{\times}$;
  2. (2)
    $\psi_{a}\in Aut(B^{\times})$ for every $a\in A^{\times}$.
Proof.
Consider an element $a\in A^{\times}$. By the definition of $\phi$, $\phi_{a}$ is a homomorphism. So we only need to show that $\phi_{a}$ is invertible. Since a homomorphism preserves the identity, $\phi(1_{A})=\phi_{1_{A}}$ is the identity endomorphism of $B$. Hence, by the properties of $\phi$,
$\phi_{a}\circ\phi_{a^{-1}}=\phi(a)\circ\phi({a^{-1}})=\phi(aa^{-1})=\phi(1_{A})=\phi(a^{-1}a)=\phi_{a^{-1}}\circ\phi_{a}.$
Thus, $\phi_{a}$ is invertible, and therefore $\phi_{a}\in Aut(B)$. This proves (1).
(2) Since a homomorphism preserves units and the restriction of a homomorphism is a homomorphism, we have that $\psi_{a}\colon B^{\times}\longrightarrow B^{\times}$ is a homomorphism for every $a\in A^{\times}$. Then an argument similar to that of (1) shows that $\psi_{a}\in Aut(B^{\times})$ for every $a\in A^{\times}$. ∎
Remark 4.7.
If $A$ and $B$ are monoids and $\phi\colon A\longrightarrow End(B)$ ($a\mapsto\phi_{a}$, $\phi_{a}\colon B\longrightarrow B$ is a homomorphism) is a homomorphism, then we have a homomorphism $\psi\colon A^{\times}\longrightarrow Aut(B^{\times})$ ($a\mapsto\psi_{a}$, where $\psi_{a}$ is the restriction of $\phi_{a}$ onto $B^{\times}$) by Lemma 4.6. Furthermore, the semi-direct product (of groups) $B^{\times}\rtimes_{\psi}A^{\times}$ with respect to $\psi$ is defined with operation inherited from that of the semi-direct product (of monoids) $B\rtimes_{\phi}A$.
Proposition 4.8.
Let $A$ and $B$ be monoids and $\phi\colon A\longrightarrow End(B)$ ($a\mapsto\phi_{a}$, $\phi_{a}\colon B\longrightarrow B$ is a homomorphism) be a homomorphism. Let $\psi\colon A^{\times}\longrightarrow Aut(B^{\times})$ be the homomorphism defined by $a\mapsto\psi_{a}$, where $\psi_{a}$ is the restriction of $\phi_{a}$ onto $A^{\times}$, then
$(B\rtimes_{\phi}A)^{\times}=B^{\times}\rtimes_{\psi}A^{\times}.$
Proof.
By the definition of $\psi$, the operation of $B^{\times}\rtimes_{\psi}A^{\times}$ is just the operation of $B\rtimes_{\phi}A$ restricted into the set $A^{\times}\bigtimes B^{\times}$. Thus, $B^{\times}\rtimes_{\psi}A^{\times}\subseteq(B\rtimes_{\phi}A)^{\times}$, and hence $B^{\times}\rtimes_{\psi}A^{\times}$ is a subgroup of $(B\rtimes_{\phi}A)^{\times}$. For the other inclusion, consider $(a,b)\in(B\rtimes_{\phi}A)^{\times}$. Then there exists $(a_{1},b_{1})\in(B\rtimes_{\phi}A)^{\times}$ such that
$(a,b)(a_{1},b_{1})=(aa_{1},b\phi_{a}(b_{1}))=(1_{A},1_{B})=(a_{1}a,b_{1}\phi_{a_{1}}(b))=(a_{1},b_{1})(a,b).$
Therefore, we have the following two equalities
$aa_{1}=a_{1}a=1_{A},\text{ and }$
$b\phi_{a}(b_{1})=b_{1}\phi_{a_{1}}(b)=1_{B}.$
The first equality implies that $a\in A^{\times}$ and $a_{1}=a^{-1}$, while the other implies $b$ is right invertible. Also, applying the automorphism $\phi_{a}$ to the equality $b_{1}\phi_{a_{1}}(b)=1_{B}$ yields since $a_{1}=a^{-1}$,
$\phi_{a}(b_{1})b=\phi_{a}(b_{1})\phi_{1_{A}}(b)=\phi_{a}(b_{1})\phi_{aa^{-1}}(b)=\phi_{a}(b_{1})\phi_{a}(\phi_{a_{1}}(b))=\phi_{a}(b_{1}\phi_{a_{1}}(b))=\phi_{a}(1_{B})=1_{B},$
hence $\phi_{a}(b_{1})b=1_{B}$, and $b$ is a left invertible. Thus, $b\in B^{\times}$ (being left and right invertible). Therefore, $(a,b)\in B^{\times}\rtimes_{\psi}A^{\times}$. ∎
Proposition 4.9.
Fix $2\leq k\leq n$. Let $\mathcal{L}^{n}_{k}$ denote the set of vector-polynomials of the form $(k:u;f)$, where $f\in R_{[k-1]}$ and $u\in R_{[k-1]}^{\times}$. Then
  1. (1)
    $\mathcal{L}^{n}_{k}$ is a subgroup of the group $TR_{n}$;
  2. (2)
    $\mathcal{L}^{n}_{k}\cong R_{[k-1]}\rtimes R_{[k-1]}^{\times}$, that is, $\mathcal{L}^{n}_{k}$ is the semi-direct product of additive group of the polynomial ring in $k-1$ variables by its group of units;
  3. (3)
    $\mathcal{L}^{n}_{k}\lhd TR_{n}$ if and only if $n=k$. In this case,
    $TR_{n}\cong\mathcal{L}^{n}_{n}\rtimes TR_{n-1}.$
Proof.
(1) Let $\mathcal{ML}^{n}_{k}$ be the monoid defined in Lemma 4.1 containing all vector-polynomials of the form $(k:u;f)$, where $f\in R_{[k-1]}$ and $u\in UV(R_{[k-1]})$. Then, by Theorem 3.13
${\mathcal{ML}^{n}_{k}}^{\times}=\mathcal{ML}^{n}_{k}\cap\mathcal{MT}_{n}^{\times}=\mathcal{ML}^{n}_{k}\cap TR_{n}=\mathcal{L}^{n}_{k}.$
(2) By Lemma 4.1, $\mathcal{ML}^{n}_{k}\cong R_{[k-1]}\rtimes UV(R_{[k-1]})$ (semi-direct product of monoids). But, by definition, $UV(R_{[k-1]})^{\times}=R_{[k-1]}^{\times}$; and $(R_{[k-1]},+)^{\times}=R_{[k-1]}$. Thus, the semi-direct product $R_{[k-1]}\rtimes R_{[k-1]}^{\times}$ is defined by Remark 4.7. Now,
[…]
(3) Assume that $k=n$. Then, by Theorem 3.3, Remark 4.7 and Proposition 4.8, we have
$TR_{n}=\mathcal{MT}_{n}^{\times}\cong(\mathcal{ML}^{n}_{n}\rtimes\mathcal{MT}_{n-1})\!^{\times}=\mathcal{L}^{n}_{n}\rtimes TR_{n-1}.$
This also, shows that $\mathcal{L}^{n}_{n}\lhd TR_{n}$. Now, assume that $2\leq k<n$ and let $(k:1;x_{k-1})\in\mathcal{L}^{n}_{k}$. So, if $(k+1:1;x_{k})\in TR_{n}\setminus\mathcal{L}^{n}_{k}$, $(k+1:1;x_{k})^{-1}=(k+1:1;-x_{k})$ by Lemma 3.11. Then, we have by direct calculations (Definition 2.6),
$(k+1:1;x_{k})(k:1;x_{k-1})(k+1:1;-x_{k})$ $=(x_{1},\ldots,x_{k-1},x_{k-1}+x_{k},x_{k-1}+x_{k+1},x_{k+2},\ldots,x_{n})$
$=(k:1;x_{k-1})(k+1:1;x_{k-1})\notin\mathcal{L}^{n}_{k}.\qed$
As a result of Proposition 4.9, we are able to decompose the group $TR_{n}$ into an iterated semi-direct product of $\mathcal{L}^{n}_{n},\ldots,\mathcal{L}^{2}_{2}$, and $TR_{1}=Aut_{R}(R[x])$. Alternatively, we can replace $\mathcal{L}^{k}_{k}$ by $R_{[k-1]}]\rtimes R_{[k-1]}^{\times}$ ($k=2,\ldots,n$) to obtain another equivalent decomposition.
Theorem 4.10.
Let $n>1$ and let $TR_{n}$ be the triangular group. Then
  1. (1)
    $TR_{n}\cong\mathcal{L}^{n}_{n}\rtimes\cdots\rtimes\mathcal{L}^{2}_{2}\rtimes Aut_{R}(R[x])$;
  2. (2)
    $TR_{n}\cong(R_{[n-1]}\rtimes R_{[n-1]}^{\times})\rtimes\cdots\rtimes(R{[x_{1}]}\rtimes R{[x_{1}]}^{\times})\rtimes Aut_{R}(R[x])$.
Moreover, $TR_{n}$ is solvable if and only if $Aut_{R}(R[x])$ is solvable. In particular, if $R$ is a domain, $TR_{n}$ is solvable, and $TR_{n}\cong(R_{[n-1]}]\rtimes R^{\times})\rtimes\cdots\rtimes(R{[x_{1}]}\rtimes R^{\times})\rtimes(R\rtimes R^{\times})$.
Proof.
By Proposition 4.9, $TR_{n}\cong\mathcal{L}^{n}_{n}\rtimes TR_{n-1}$. Then iterated application of this relation yields the first expression. Again, by Proposition 4.9, we have $\mathcal{L}^{i}_{i}\cong R_{[i-1]}\rtimes R_{[i-1]}^{\times}$ for $i=2,\ldots,n$. Combining these results yields the second expression. Moreover, it is well known that Abelian groups are solvable and extensions of solvable groups are solvable. Therefore, by the second expression, $TR_{n}$ is solvable if and only if $Aut_{R}(R[x])$ is solvable. In particular, if $R$ is a domain, then $Aut_{R}(R[x])\cong R\rtimes R^{\times}$ and for each $i$, $R_{[i]}^{\times}=R^{\times}$. Thus, $TR_{n}$ is solvable, and the claim follows. ∎
Remark 4.11.
  1. (1)
    In algebraic geometry, the Jonquières (classical) triangular group $KTR_{n}$ is defined to be the group of all vector-polynomials of the form
    $\vec{f}=(a_{1}x_{1}+b_{0},f_{1}(x_{1})+a_{2}x_{2},\ldots,f_{n-1}(x_{1},\ldots,x_{n-1})+a_{n}x_{n}),$ (8)
    where $a_{1},\ldots,a_{n}\in R^{\times}$ and $b\in R$ (see [25]). In general, we have the inclusions:
    $KTR_{n}\subseteq TR_{n}\subseteq\mathcal{MT}_{n}.$ (9)
    However, by Theorem 4.10 and Theorem 4.5, one can easily see that
    $KTR_{n}=TR_{n}=\mathcal{MT}_{n}$
    if and only if:
    1. (a)
      $UV(R_{[i]})=R_{[i]}^{\times}=R^{\times}$.
    2. (b)
      $\mathcal{MP}(R)=Aut_{R}(R[x])=\{ax+b\colon a\in R^{\times}\}$.
    In the following examples, we demonstrate how the validity of the previous two conditions determines the nature of the inclusions of relation (9).
    • When $R$ is a $D$-ring (see for example [13]), $R$ satisfies condition (1a) by the definition of $D$-rings. Then $KTR_{n}=TR_{n}$ (note that it is enough here to have $R_{[i]}^{\times}=R^{\times}$).
    • If $R$ is a non-algebraically closed field (for example, real numbers and finite fields),
      $KTR_{n}=TR_{n}\subsetneq\mathcal{MT}_{n}.$
      In this case, it is clear that $R_{[i]}^{\times}=R^{\times}$ (for each $i$) and $Aut_{R}(R[x])$ consists of all linear polynomials. Also, since $R$ is not an algebraically closed field, there is an irreducible polynomial $f$ of degree $\geq 2.$ Evidently, $f\in UV(R_{[i]})\setminus R_{[i]}^{\times}$.
    • When $R$ is not reduced (i.e. containing a non-zero nilpotent), we have
      $KTR_{n}\subsetneq TR_{n}\subseteq\mathcal{MT}_{n}.$
      Here, we have by Remark 2.4, $1+rx_{1}\in R_{[i]}^{\times}\neq R^{\times}$ for any non-zero nilpotent $r$. In particular, if $R$ is a finite local ring that is not a field,
      $KTR_{n}\subsetneq TR_{n}\subsetneq\mathcal{MT}_{n},$
      since in this case we also have $1+(x^{q}-x)\in UV(R[x])\setminus R[x]^{\times}$.
  2. (2)
    In Equation (8), if $a_{1}=\cdots=a_{n}=1$, we get the so called unitriangular group of automorphisms. In [2], a decomposition of the unitriangular group on fields of characteristic zero into an iterated semi-product of Abelian groups is given.
For algebraically closed fields we have the following result.
Proposition 4.12.
Let $R$ be an algebraically closed field.
  1. (1)
    If $R$ of characteristic zero, then
    $KTR_{n}=TR_{n}=\mathcal{MT}_{n}.$
  2. (2)
    If $R$ of characteristic $p\neq 0$, then
    $KTR_{n}=TR_{n}\subsetneq\mathcal{MT}_{n}.$
Proof.
First assume that $R$ is any algebraically closed field. Then $UV(R_{[i]})=R_{[i]}^{\times}=R^{\times}$ since every non-constant polynomial has a root over an algebraically closed field. Also, since $R$ is a field, $Aut_{R}[x]=\{ax+b\colon a\in R^{\times}\}$. Thus, $KTR_{n}=TR_{n}$, by Remark 4.11.
Now, assume that $R$ is an algebraically closed field of characteristic zero. Then $\mathcal{MP}(R)=Aut_{R}(R[x])$ since $R$ has no non-linear permutation polynomial by [4, Theororem 6]. This together with the first part of the proof implies $TR_{n}=\mathcal{MT}_{n}$ by Remark 4.11. This proves (1).
For the case where $R$ is an algebraically closed field of characteristic $p\neq 0$, we have that $x^{p}\in\mathcal{MP}(R)\setminus Aut_{R}(R[x])$ by Remark  3.5 (3). Thus, $TR_{n}\subsetneq\mathcal{MT}_{n}$ by Remark 4.11. ∎
Remark 4.13.
Unlike solvability, the decomposition of Theorem 4.10 can not determine whether the group $TR_{n}$ is nilpotent or not. However, in the following we notice for a large class of rings that $TR_{n}$ is not nilpotent.
  1. (1)
    Let $n>1$ and $R$ is an infinite field. Then $TR_{n}$ is not nilpotent.
  2. (2)
    Let $n>2$ and $R$ be a commutative ring with Jacobson radical $J(R)$ such that $R/J(R)$ is the finite field of $q$ elements $\mathbb{F}_{q}$.
To show these assertions, it is enough to find a subgroup of $TR_{n}$, which is not nilpotent.
Indeed the first assertion follows from the fact that the unitriangular group (defined in  Remark 4.11) is not nilpotent (see [2, Theorem 2.1]). To prove the second assertion, we consider first the case $R=\mathbb{F}_{q}$. But, then the unitrianguar group is not nilpotent (see [12, Theorem 3]). For the case $J(R)\neq\{0\}$, we note that there exists a natural epimorphism $\pi\colon R\longrightarrow R/J(R)$ given by $\pi(a)=a\mod J(R)$. Evidently $\pi$ can be extended in to an epimorphism of monoids $\tilde{\pi}\colon R_{[n]}^{n}\longrightarrow({R/J(R)})_{[n]}^{n}$, and one easily see that $\tilde{\pi}(UT_{n}(R))=UT_{n}(R/J(R))=UT_{n}(\mathbb{F}_{q})$, where $UT_{n}(A)$ stands for the unitriangular group over the ring $A$. Hence $UT_{n}(R)$ is not nilpotent since $\tilde{\pi}(UT_{n}(R))$ is not nilpotent. Thus, $TR_{n}$ is not nilpotent.
4.3. The structure of the induced permutations group $\pi_{n}(\mathcal{MT}_{n})$
Recall that $\pi_{n}$ stands for the natural homomorphism that assigns for each $\vec{f}\in R_{[n]}^{n}$ its induced vector-function $\vec{F}\colon R^{n}\longrightarrow R^{n}$.
In this subsection, we explore the structure of the group $\pi_{n}(\mathcal{MT}_{n})$ of permutations of $R^{n}$ induced by the elements of the triangular monoid $\mathcal{MT}_{n}$. We see that this group can be decomposed into an iterated semi-direct product of $n$ groups in a similar manner to what we have already seen for the group $TR_{n}$.
Proposition 4.14.
Let $R$ be a finite commutative ring and fix $2\leq k\leq n$. Let $\mathcal{ML}^{n}_{k}$ be the monoid of vector-permutation polynomials of the form $(k:u;f)$, where $f\in R_{[k-1]}$ and $u\in UV(R_{[k-1]})$. Then
  1. (1)
    $\pi_{n}(\mathcal{ML}^{n}_{k})$ is a subgroup of the group $\pi_{n}(\mathcal{MT}_{n})$;
  2. (2)
    $\pi_{n}(\mathcal{ML}^{n}_{k})\cong\mathcal{F}(R^{k-1})\rtimes\mathcal{F}(R^{k-1})^{\times}$, that is, $\mathcal{ML}^{n}_{k}$ is the semi-direct product of the additive group of the ring of polynomial functions (on $R$) in $k-1$ variables by its group of units;
  3. (3)
    $\pi_{n}(\mathcal{ML}^{n}_{k})\lhd\pi_{n}(\mathcal{MT}_{n})$ if and only if $n=k$. In this case,
    $\pi_{n}(\mathcal{MT}_{n})\cong\pi_{n}(\mathcal{ML}^{n}_{n})\rtimes\pi_{n}(\mathcal{MT}_{n-1}).$
Proof.
Let $k$ be fixed. (1) Obvious.
(2) By Lemma 4.1, $\mathcal{ML}^{n}_{k}\cong R_{[k-1]}\rtimes UV(R_{[k-1]})$. Then it is easy to see that the sets $A=\{(k:1;f):f\in R_{[k-1]}\}$ and $B=\{(k:u;0):u\in UV(R_{[k-1]}\}$ are isomorphic to $R_{[k-1]}$ and $UV(R_{k-1})$ respectively. Further, $\pi_{n}(A)$ and $\pi_{n}(B)$ are subgroups of the finite group $\pi_{n}(\mathcal{ML}^{n}_{k})$ (being non empty and closed with respect to composition).
Now, we claim that $\pi_{n}(A)\cong\mathcal{F}(R^{k-1})$ and $\pi_{n}(B)\cong\mathcal{F}(R^{k-1})^{\times}$. To show this claim, let $\pi\colon R_{[k-1]}\longrightarrow\mathcal{F}(R^{k-1})$ be the natural epimorphism that maps every polynomial to its induced function on $R$. Then define a map $\Psi\colon\pi_{n}(A)\longrightarrow\mathcal{F}(R^{k-1})$ by $\Psi(\vec{F})=\pi(f)$, where $f\in R_{[k-1]}$ such that $\vec{F}=\pi_{n}((k:1;f))$. By Lemmas 2.13 and 3.12, $\Psi$ is well defined and injective. Let $\vec{G}\in\pi_{n}(A)$. Then $\vec{G}=\pi_{n}((k:1;g))$ for some $g\in R_{[k-1]}$. Consider,
$\Psi(\vec{F}\vec{G})$ $=\Psi(\pi_{n}((k:1;f))\circ\pi_{n}((k:1;g)))=\Psi(\pi_{n}((k:1;f)(k:1;g)))$
[…]
$=\pi(f)+\pi(g)=\Psi(\vec{F})+\Psi(\vec{G}).$
This shows that $\Psi$ is a homomorphism. By the definitions of the set $A$, and the maps $\Psi$, $\pi_{n}$ and $\pi$, we have
$\Psi(\pi_{n}(A))=\pi(R_{[k-1]})=\mathcal{F}(R^{k-1}),$
whence $\Psi$ is surjective. Therefore, $\pi_{n}(A)\cong\mathcal{F}(R^{k-1})$. Similarly, one can show that $\pi_{n}(B)\cong\mathcal{F}(R^{k-1})^{\times}$.
Now consider the map $\Phi\colon\pi_{n}(\mathcal{ML}^{n}_{k})\longrightarrow\pi_{n}(B),(\pi_{n}((k:u;f))\mapsto\pi_{n}((k:u;0)))$. It is obvious that $\Phi$ is surjective. So, if $\pi_{n}((k:u;f)),\pi_{n}((k:v;g))\in\pi_{n}(\mathcal{ML}^{n}_{k})$, we have
$\Phi(\pi_{n}((k:u;f))\circ\pi_{n}((k:v;g)))$ $=\Psi(\pi_{n}((k:u;f)(k:v;g)))$
$=\Phi(\pi_{n}((k:uv;f+ug))=(\pi_{n}((k:uv;0))$
$=\pi_{n}((k:u;0)(k:v;0))=\pi_{n}((k:u;0)\circ\pi_{n}((k:v;0))$
$=\Phi(\pi_{n}((k:u;f)))\circ\Phi(\pi_{n}((k:v;g))).$
Thus, $\Phi$ is an epimorphism. Now, if $j\colon\pi_{n}(B)\longrightarrow\pi_{n}(\mathcal{ML}^{n}_{k})$ is the inclusion, then evidently, $\Phi j=id_{\pi_{n}(B)}$. Therefore, the following extension splits
[…]
Thus,
$\pi_{n}(\mathcal{ML}^{n}_{k})=\pi_{n}(A)\rtimes\pi_{n}(B)\cong\mathcal{F}(R^{k-1})\rtimes\mathcal{F}(R^{k-1})^{\times}.$
(3) Let $\vec{F}\in\pi_{n}(\mathcal{MT}_{n})$. Then $\vec{F}=\pi_{n}(\vec{f})$ for some $\vec{f}=(f_{1},\ldots,f_{n-1},f_{n})$ with $f_{i}\in R_{[i]}$ for $i=1,\ldots,n$ (by the definition of the triangular monoid $\mathcal{MT}_{n}$). Define $P\colon\pi_{n}(\mathcal{MT}_{n})\longrightarrow\pi_{n-1}(\mathcal{MT}_{n-1})$ by letting $P(\vec{F})=\pi_{n-1}((f_{1},\ldots,f_{n-1}))$. Clearly, $P$ is an epimorphism. Also, the map $j\colon\pi_{n-1}(\mathcal{MT}_{n-1})\longrightarrow\pi_{n}(\mathcal{MT}_{n})$, $(\pi_{n-1}((f_{1},\ldots,f_{n-1})))\mapsto\pi_{n}((f_{1},\ldots,f_{n-1},x_{n}))$ is obviously a monomorphism. Further, $Pj=id_{\pi_{n-1}(\mathcal{MT}_{n-1})}$. Hence, the following extension splits
[…]
Therefore, $\pi_{n}(\mathcal{MT}_{n})\cong\pi_{n}(\mathcal{ML}^{n}_{k})\rtimes\pi_{n-1}(\mathcal{MT}_{n-1}),\text{ and }\pi_{n}(\mathcal{ML}^{n}_{k})\lhd\pi_{n}(\mathcal{MT}_{n}).$
Now, assume that $2\leq k<n$. Then, we adopt the argument given in the last part of the proof of Proposition 4.9 (but with applying the homomorphism $\pi_{n}$). So, if $\pi_{n}((k+1:1;x_{k}))\in\pi_{n}(TR_{n})\subseteq\pi_{n}(\mathcal{MT}_{n})$, $(\pi_{n}((k+1:1;x_{k})))^{-1}=\pi_{n}((k+1:1;x_{k})^{-1})=\pi_{n}((k+1:1;-x_{k}))$. Consider the element $\pi_{n}(k:1;x_{k-1})\in\pi_{n}(\mathcal{ML}^{n}_{k})$, we have since $\pi_{n}$ is a homomorphism,
$\pi_{n}((k+1:1;x_{k})(k:1;$ $x_{k-1})(k+1:1;-x_{k}))$
$=\pi_{n}((x_{1},\ldots,x_{k-1},x_{k-1}+x_{k},x_{k-1}+x_{k+1},x_{k+2},\ldots,x_{n}))$
$=\pi_{n}((k:1;x_{k-1}))\pi_{n}((k+1:1;x_{k-1})).$
Thus, since $\pi_{n}((k:1;x_{k-1}))\in\pi_{n}(\mathcal{ML}^{n}_{k})$, $\pi_{n}((k:1;x_{k-1}))\pi_{n}((k+1:1;x_{k-1}))\notin\pi_{n}(\mathcal{ML}^{n}_{k})$ if and only if $\pi_{n}((k+1:1;x_{k-1}))\notin\pi_{n}(\mathcal{ML}^{n}_{k})$. Let us assume that $\pi_{n}((k+1:1;x_{k-1}))\in\pi_{n}(\mathcal{ML}^{n}_{k})$. By the definition of $\pi_{n}(\mathcal{ML}^{n}_{k})$, there exist $\vec{f}\in\ \mathcal{ML}^{n}_{k}$ such that $\pi_{n}(\vec{f})=\pi_{n}((k+1:1;x_{k-1}))$. Hence, there exist $f\in R_{[k-1]}$ and $u\in UV(R_{[k-1]})$ such that $\vec{f}=(k:u;f)$ by the definition of the monoid $\mathcal{ML}^{n}_{k}$. But, then by Lemma 3.12, $(k:u;f)\operatorname{{}\triangleq{}}(k+1:1;x_{k-1})$ on $R^{n}$, which implies that $x_{k+1}\operatorname{{}\triangleq{}}x_{k+1}+x_{k-1}$ on $R$ by Lemma 2.13. Thus, $x_{k-1}\operatorname{{}\triangleq{}}0$ on $R$, that is the polynomial $x_{k-1}$ maps $R^{n}$ to the zero, which is impossible since $R$ contains $1$. Therefore, $\pi_{n}((k+1:1;x_{k-1}))\notin\pi_{n}(\mathcal{ML}^{n}_{k})$ and $\pi_{n}(\mathcal{ML}^{n}_{k})$ is not normal subgroup of $\pi_{n}(\mathcal{MT}_{n})$. ∎
Similar to Theorem 4.10, we view the group $\pi_{n}(\mathcal{MT}_{n})$ as iterated semi-direct products of groups. Also, we obtain a counting formula for its order.
Theorem 4.15.
Let $n>1$ and let $R$ be a finite commutative ring. Then
  1. (1)
    $\pi_{n}(\mathcal{MT}_{n})\cong\pi_{n}(\mathcal{ML}^{n}_{n})\rtimes\cdots\rtimes\pi_{2}(\mathcal{ML}^{2}_{2})\rtimes\mathcal{P}(R)$;
  2. (2)
    $\pi_{n}(\mathcal{MT}_{n})\cong(\mathcal{F}(R^{n-1})\rtimes\mathcal{F}(R^{n-1})^{\times})\rtimes\cdots\rtimes(\mathcal{F}(R)\rtimes\mathcal{F}(R)^{\times})\rtimes\mathcal{P}(R)$;
  3. (3)
    $|\pi_{n}(\mathcal{MT}_{n})|=|\mathcal{P}(R)|\times\prod_{i=1}^{n-1}|\mathcal{F}(R^{i})||\mathcal{F}(R^{i})^{\times}|.$
Proof.
(3) follows from (2). (2) follows from (1) and Proposition 4.14-(2). By Proposition 4.14-(3), $\pi_{n}(\mathcal{MT}_{n})\cong\pi_{n}(\mathcal{ML}^{n}_{n})\rtimes\pi_{n}(\mathcal{MT}_{n-1})$. Then one can use this relation and induction to prove (1). We note to the reader that $\mathcal{MT}_{1}=\mathcal{MP}(R)$ and $\pi_{1}(\mathcal{MP}(R))=\mathcal{P}(R)$. ∎
The second decomposition of the previous theorem will allow obtaining more information about the group $\pi_{n}(\mathcal{MT}_{n})$. However, without loss of generality, we restrict ourselves to finite local rings. We will need some facts about polynomial functions over finite local rings. In the proof of the following proposition, we use the fact that the order of a finite local commutative ring is a power of $q$, where $q$ is the number of elements of its residue field $\mathbb{F}_{q}$.
Proposition 4.16.
Let $R$ be a finite local ring and $k\geq 1$. Then the group $(\mathcal{F}(R^{k}),``+")$ is a $p$-group.
Proof.
Let $F\in\mathcal{F}(R^{k})$, then there exists a polynomial $f_{F}\in R_{[k]}$ of minimal degree representing $F$. Since $\mathcal{F}(R^{k})$ is finite, there exists a positive number $n$ such that $\deg f_{F}\leq n$ for every $F\in\mathcal{F}(R^{k})$. Let
$A=\{g\in R_{[k]}\colon\deg g\leq n\}.$
Then $A$ is a group with respect to addition. Further, $|A|=|R|^{1+\sum_{i=1}^{k}l_{i}}=q^{({1+\sum_{i=1}^{k}l_{i}})}$, where $l_{i}$ is the number of monomials of degree $i$. Hence, $A$ is a $p$-group since $q$ is a power of a prime $p$ (say $q=p^{r}$ for some positive integer). Now, define a map $\psi\colon A\longrightarrow\mathcal{F}(R^{k})$ as follows: if $F$ is the function induced by $f$, we let $\psi(f)=F$. Clearly, $\psi$ is a homomorphism. It is also surjective since $f_{F}\in A$ for every $F\in\mathcal{F}(R^{k})$. Thus, by the First Isomorphism Theorem of groups,
[…]
But, $|\ker\psi|=p^{m}$ for some non-negative integer $m$ (being a subgroup of a $p$-group). Therefore, $\mathcal{F}(R^{k})$ is a $p$-group. ∎
Theorem 4.17.
Let $R$ be a finite local ring and $k\geq 1$. Then
$\frac{|\mathcal{F}(R^{k})^{\times}|}{|\mathcal{F}(R^{k})|}=\frac{(q-1)^{q^{k}}}{q^{q^{k}}}.$
Proof.
Set $A=R/M\cong\mathbb{F}_{q}$. Since every function $F\colon\mathbb{F}_{q}^{k}\longrightarrow\mathbb{F}_{q}$ is a polynomial function, we have $|\mathcal{F}(A^{k})|=q^{{q^{k}}}$. Also, since $|A^{\times}|=(q-1)$ we have $|\mathcal{F}(R^{k})^{\times}|=(q-1)^{q^{k}}$. Since polynomial functions preserve congruences modulo $M$, there exists a natural ring epimorphism $\lambda\colon\mathcal{F}(R^{k})\longrightarrow\mathcal{F}(A^{k})$. Thus,
$|\mathcal{F}(R^{k})|=|\ker\lambda||\mathcal{F}(A^{k})|=q^{q^{k}}|\ker\lambda|.$
Now, an element $a\in R$ is a unit in $R$ if and only if $a+M$ is a unit in $A$. Therefore, $F\in\mathcal{F}(R^{k})^{\times}$ if and only if $\lambda(F)\in\mathcal{F}(A^{k})^{\times}$. This yields, $\mathcal{F}(R^{k})^{\times}=\lambda^{-1}(\mathcal{F}(A^{k})^{\times})$, whence
$|\mathcal{F}(R^{k})^{\times}|=|\ker\lambda||\mathcal{F}(A^{k})^{\times}|=(q-1)^{q^{k}}|\ker\lambda|.$
Therefore,
$\frac{|\mathcal{F}(R^{k})^{\times}|}{|\mathcal{F}(R^{k})|}=\frac{(q-1)^{q^{k}}|\ker\lambda|}{q^{q^{k}}|\ker\lambda|}=\frac{(q-1)^{q^{k}}}{q^{q^{k}}}.\qed$
Corollary 4.18.
Let $R$ be a finite local ring that is not a field with residue field $\mathbb{F}_{2}$. Then $\mathcal{F}(R^{k})^{\times}$ is a $2$-group.
Proof.
By Theorem 4.17,
$|\mathcal{F}(R^{k})^{\times}|=\frac{(2-1)^{2^{k}}}{2^{2^{k}}}|\mathcal{F}(R^{k})|=\frac{|\mathcal{F}(R^{k})|}{2^{2^{k}}}.$
Since $R\neq\mathbb{F}_{2}$, the proof of Theorem 4.17 shows that $|\mathcal{F}(R^{k})|>|\mathcal{F}(\mathbb{F}_{2}^{k})|={2^{2^{k}}}$. The result follows from Proposition 4.16. ∎
Remark 4.19.
Similar to Theorem 4.17, but in one variable, Jiang [10] has obtained the following relation between the number of polynomial functions and the number of polynomial permutations on a finite local ring $R$ which is not a field:
$\frac{|\mathcal{F}(R)|}{|\mathcal{P}(R)|}=\frac{q!(q-1)^{q}}{q^{2q}}.$ (10)
Lemma 4.20.
[8, Theorem 4.1] Let $R$ be a finite local commutative ring with a residue field of $q$ elements $\mathbb{F}_{q}$. Let $\mathcal{P}(R)$ be the group of polynomial permutations. Then
  1. (1)
    $\mathcal{P}(R)$ is solvable if and only if $q\leq 4$;
  2. (2)
    $\mathcal{P}(R)$ is nilpotent if and only if $q=2$;
  3. (3)
    $\mathcal{P}(R)$ is abelian if and only if $R=\mathbb{F}_{2}$.
With these preparations, we are ready to present our main theorem of this subsection.
Theorem 4.21.
Let $n\geq 1$ and let $R$ be a finite local commutative ring with a residue field of $q$ elements $\mathbb{F}_{q}$. Let $\pi_{n}(\mathcal{MT}_{n})$ be the group of permutations on $R^{n}$ induced by the monoid $\mathcal{MT}_{n}$. Then
  1. (1)
    $\pi_{n}(\mathcal{MT}_{n})$ is solvable if and only if $q\leq 4$;
  2. (2)
    $\pi_{n}(\mathcal{MT}_{n})$ is nilpotent if and only if $q=2$;
  3. (3)
    $\pi_{n}(\mathcal{MT}_{n})$ is abelian if and only if $n=1$ and $R=\mathbb{F}_{2}$
Proof.
In the light of Lemma 4.20 we may assume that $n>1$ since $\pi_{1}(\mathcal{MT}_{1})=\mathcal{P}(R)$.
By Theorem 4.15,
$\pi_{n}(\mathcal{MT}_{n})\cong(\mathcal{F}(R^{n-1})\rtimes\mathcal{F}(R^{n-1})^{\times})\rtimes\cdots\rtimes(\mathcal{F}(R)\rtimes\mathcal{F}(R)^{\times})\rtimes\mathcal{P}(R).$
Now, $\mathcal{F}(R^{i})$ and $\mathcal{F}(R^{i})^{\times}$ are solvable groups (being abelian groups). Also, extensions of solvable groups are solvable. Therefore $\pi_{n}(\mathcal{MT}_{n})$ is solvable if and only if $\mathcal{P}(R)$ is solvable if and only if $q\leq 4$ by Lemma 4.20. This proves (1).
(2) Since every subgroup of a nilpotent group is nilpotent, we need only to consider the case $q=2$ by Lemma 4.20. Assume that $q=2$ and without loss of generality we consider the case $R\neq\mathbb{F}_{q}$. We claim that $\pi_{n}(\mathcal{MT}_{n})$ is a $2$-group, hence it is nilpotent since $p$-groups are nilpotents. To show the claim, we notice that by means of the above the decomposition of $\pi_{n}(\mathcal{MT}_{n})$, $\pi_{n}(\mathcal{MT}_{n})$ is a $2$-group if and only if $\mathcal{P}(R)$, $\mathcal{F}(R^{i})$ and $\mathcal{F}(R^{i})^{\times}$ are $2$-groups for $i=1,\ldots,n-1$. Proposition 4.16 shows that $\mathcal{F}(R^{i})$ is a $2$-group, while $\mathcal{F}(R^{i})^{\times}$ is a $2$-group by Corollary 4.18 for $i=1,\ldots,n-1$. Finally, using Equation (10), we see that $\mathcal{P}(R)$ is a $2$-group.
To prove (3), we need only to consider the case $R=\mathbb{F}_{2}$ and $n=2$ since $\pi_{2}(\mathcal{MT}_{2})$ is embedded in $\pi_{n}(\mathcal{MT}_{n})$ for every $n>2$ by Remark 3.19. The routine calculations show that
$\pi_{2}((x_{1},x_{1}+x_{2}))\circ\pi_{2}((x_{1}+1,x_{1}+x_{2}))\neq\pi_{2}((x_{1}+1,x_{1}+x_{2}))\circ\pi_{2}((x_{1},x_{1}+x_{2})).\qed$
5. A connection to the group of polynomial permutations of the ring dual numbers of several variables
Given a commutative ring $R$ with unity, we define the ring of dual numbers of $n\geq 1$ variables over $R$ to be the quotient ring $R_{[n]}/I$, where $I$ is the ideal generated by the set $\{x_{i}x_{j}\mid i,j\in\{1,\ldots,n\}\}$. This quotient ring contains an isomorphic copy of $R$. Also, this ring can be viewed as the free commutative $R$-algebra ${R}[\operatorname{\alpha}_{1},\ldots,\operatorname{\alpha}_{n}]$ with basis $1,\operatorname{\alpha}_{1},\ldots,\operatorname{\alpha}_{n}$ such that $\operatorname{\alpha}_{i}$ stands for $x_{i}+I$ and $\operatorname{\alpha}_{i}\operatorname{\alpha}_{j}=0$ for $i,j=1,\ldots,n$.
To simplify notation we shall use $R_{n}$ instead of ${R}[\operatorname{\alpha}_{1},\ldots,\operatorname{\alpha}_{n}]$ (or $R_{[n]}/I$). We observe here that every element $r\in R_{n}$ can be written uniquely as
$r=r_{0}+\sum\limits_{i=1}^{n}r_{i}\operatorname{\alpha}_{i}\text{ for some }r_{0},\ldots,r_{n}\in R.$
Because of this every polynomial $g\in R_{n}[x]$ has a unique representation
$g=g_{0}+\sum\limits_{i=1}^{n}g_{i}\operatorname{\alpha}_{i},\text{ for some }g_{0},g_{1},\ldots,g_{n}\in R[x].$
Also, it is evident that $r_{0}+\sum_{i=1}^{n}r_{i}\mapsto(r_{0},\ldots,r_{n})$ establishes a one-to-one correspondence between $R_{n}$ and $R^{n+1}$. This correspondence motivates us to relate some groups of permutations of $R_{n}$ and $R$, at least in the finite case.
Throughout this section, since we deal with two different rings, we use the notation $\mathcal{MT}_{n+1}(R)$ for the triangular mononid over the ring $R$ to avoid any confusion with the coefficient ring.
In this section, for a commutative ring $R$, we embed the monoid $\mathcal{MP}(R_{n})$ of permutation polynomials on $R_{n}$ in the triangular monoid $\mathcal{MT}_{n+1}(R)$. Moreover, this embedding leads to the embedding of the group of polynomial permutations $\mathcal{P}(R_{n})$ in the group of $\pi_{n+1}(\mathcal{MT}_{n+1}(R))$, whenever $R$ is finite.
Lemma 5.1.
Let $k>1$, $R$ be a commutative ring and $g=g_{0}+\sum\limits_{i=1}^{k}g_{i}\operatorname{\alpha}_{i}$, where $g_{0},\ldots,g_{k}\in R[x]$.
  1. (1)
    If $f\in R[x]$, then
    $f(g)=f(g_{0})+\sum\limits_{i=1}^{k}g_{i}f^{\prime}(g_{0})\operatorname{\alpha}_{i}.$
  2. (2)
    If $f=f_{0}+\sum\limits_{i=1}^{k}f_{i}\operatorname{\alpha}_{i}$, where $f_{0},\ldots,f_{k}\in R[x]$, then
    $f(g)=f_{0}(g_{0})+\sum\limits_{i=1}^{k}(g_{i}f_{0}^{\prime}(g_{0})+f_{i}(g_{0}))\operatorname{\alpha}_{i}.$
Proof.
(2) Follows from (1). To show (1), write $f(x)=\sum\limits_{j=0}^{m}a_{j}x^{j}$. Then expanding $f(g)$ with the fact that $\operatorname{\alpha}_{i}\operatorname{\alpha}_{j}=0$ for all $i,j$ yields the result. ∎
We need the following facts from [1].
Lemma 5.2.
[1, Theorem 4.1] Let $R$ be a finite ring. Let $f=f_{0}+\sum\limits_{i=1}^{k}f_{i}\operatorname{\alpha}_{i}$, where $f_{0},\ldots,f_{k}\in R[x]$. Then $f$ is a permutation polynomial on $R_{k}$ if and only if the following conditions hold:
  1. (1)
    $f_{0}$ is a permutation polynomial on $R$;
  2. (2)
    $f_{0}$ is a unit-valued polynomial on $R$.
Lemma 5.3.
[1, Corollary 3.6]
Let $f=f_{0}+\sum\limits_{i=1}^{k}f_{i}\operatorname{\alpha}_{i}$ and $g=g_{0}+\sum\limits_{i=1}^{k}g_{i}\operatorname{\alpha}_{i}$, where $f_{0},\ldots,f_{k},g_{0},\ldots,g_{k}\in R[x]$.
Then $f$ and $g$ induce the same function on $R_{k}$ if and only if the following conditions hold:
  1. (1)
    $f_{i}\operatorname{{}\triangleq{}}g_{i}$ on $R$ for $i=0,\dots,k$;
  2. (2)
    $f_{0}^{\prime}\operatorname{{}\triangleq{}}g_{0}^{\prime}$ on $R$.
We conclude this section with the following embedding result.
Proposition 5.4.
Let $n>1$ and let $R$ be a commutative ring. Then
  1. (1)
    $\mathcal{MP}(R_{n})\hookrightarrow\mathcal{MT}_{n+1}(R)$; and whenever $R$ is finite
  2. (2)
    $\mathcal{P}(R_{n})\hookrightarrow\pi_{n+1}(\mathcal{MT}_{n+1}(R))$.
Proof.
(1) Let $f=f_{0}+\sum\limits_{i=1}^{n}f_{i}\operatorname{\alpha}_{i}\in\mathcal{MP}(R_{n})$, where $f_{0},\ldots,f_{n}\in R[x]$. Then, $f_{0}$ is a permutation polynomial and $f_{0}^{\prime}$ is a unit-valued polynomial by Lemma 5.2. Define $\psi\colon\mathcal{MP}(R_{n})\longrightarrow\mathcal{MT}_{n+1}(R)$ by
$\psi(f)=(f_{0}(x_{1}),f_{1}(x_{1})+x_{2}f_{0}^{\prime}(x_{1}),\ldots,f_{n}(x_{1})+x_{n+1}f_{0}^{\prime}(x_{1})).$
Since $f_{0}$ is a permutation polynomial and $f_{0}^{\prime}$ is a unit-valued polynomial, $\psi(f)$ has the form (3) in Theorem 3.3. Thus, $\psi(f)\in\mathcal{MT}_{n+1}$ and $\psi$ is defined. Now, let $g=g_{0}+\sum\limits_{i=1}^{n}g_{i}\operatorname{\alpha}_{i}\in\mathcal{MP}(R_{n})$, where $g_{0},\ldots,g_{n}\in R[x]$. By Lemma 5.1, $f\circ g(x)=f_{0}(g_{0}(x))+\sum\limits_{i=1}^{n}(g_{i}(x)f_{0}^{\prime}(g_{0}(x))+f_{i}(g_{0}(x)))\operatorname{\alpha}_{i}$.
Thus, by the definition of $\psi$ and the definition of composition,
$\psi(f\circ g)$ $=(f_{0}(g_{0}(x_{1})),(g_{1}(x_{1})f_{0}^{\prime}(g_{0}(x_{1}))+f_{1}(g_{0}(x_{1})))+x_{2}g_{0}^{\prime}(x_{1})f_{0}^{\prime}(g_{0}(x_{1})),\ldots$
$\quad\ldots,(g_{n}(x_{1})f_{0}^{\prime}(g_{0}(x_{1}))+f_{n}(g_{0}(x_{1})))+x_{n+1}g_{0}^{\prime}(x_{1})f_{0}^{\prime}(g_{0}(x_{1})))$
$=\psi(f)\circ\psi(g).$
Thus, $\psi$ is a homomorphism. Evidently, $\psi$ is one-to-one.
(2) Define $\phi\colon\mathcal{P}(R_{n})\longrightarrow\pi_{n+1}(\mathcal{MT}_{n+1})$ by $\phi(F)=\pi_{n+1}(\psi(f))$, where $f=f_{0}+\sum\limits_{i=1}^{n}f_{i}\operatorname{\alpha}_{i}\in\mathcal{MP}(R_{n})$ is a permutation polynomial inducing $F$ on $R_{n}$, with $f_{0},\ldots,f_{n}\in R[x]$. Then, by Lemma 5.3 and Lemma 3.12, $\phi$ is well defined. Also, $\phi$ is a homomorphism as it is a composition of homomorphisms. To show $\phi$ is injective, consider $G\in\mathcal{P}(R_{n})$ such that $\phi(F)=\phi(G)$. So, there exists $g=g_{0}+\sum\limits_{i=1}^{n}g_{i}\operatorname{\alpha}_{i}\in\mathcal{MP}(R_{n})$, where $g_{0},\ldots,g_{n}\in R[x]$ such that $G$ is induced by $g$. Then $\phi(F)=\phi(G)$ implies that $\pi_{n+1}(\psi(g))=\pi_{n+1}(\psi(f))$, that is
$\pi_{n+1}((f_{0}(x_{1}),f_{1}(x_{1})+x_{2}f_{0}^{\prime}(x_{1})$ $,\ldots,f_{n}(x_{1})+x_{n+1}f_{0}^{\prime}(x_{1})))=$
$=\pi_{n+1}((g_{0}(x_{1}),g_{1}(x_{1})+x_{2}g_{0}^{\prime}(x_{1}),\ldots,g_{n}(x_{1})+x_{n+1}g_{0}^{\prime}(x_{1}))).$
By Lemma 3.12, we have
$\begin{matrix}f_{0}(x_{1})\operatorname{{}\triangleq{}}g_{0}(x_{1})\text{ on }R,\text{ and }\\ f_{i}(x_{1})+x_{i+1}f_{0}^{\prime}(x_{1})\operatorname{{}\triangleq{}}g_{i}(x_{1})+x_{i+1}g_{0}^{\prime}(x_{1})\text{ on }R\text{ for }i=1,\dots,n.\end{matrix}$
Hence, by Lemma 2.13, $f_{0}^{\prime}(x_{1})\operatorname{{}\triangleq{}}g_{0}^{\prime}(x_{1})$ and $f_{i}(x_{1})\operatorname{{}\triangleq{}}g_{i}(x_{1})$ for $i=1,\ldots,n$. Now, the conditions of Lemma 5.3 are valid, and therefore, $f$ and $g$ induce the same function on $R_{n}$, that is $F=G$. This finishes the proof. ∎
6. The tame monoid and related questions
Recall from Remark 4.11 that the classical triangular group $KTR_{n}$ consists of all invertible vector-polynomials of the form $(a_{1}x_{1}+a_{1},f_{1}(x_{1})+a_{2}x_{2},\ldots,f_{n-1}(x_{1},\ldots,x_{n-1})+ax_{n})$, where $a_{0}\in R$ and $a_{1},\ldots,a_{n}\in R^{\times}$. It is well-known that the group $KTR_{n}$ and the affine Group $Aff_{n}$ consisting of all invertible linear vector-polynomials generate the so-called tame group $Tm_{n}$. An element of $Tm_{n}$ is called tame, and the tame question asks (when $R$ is a field and $n\geq 2$) whether the group $GA_{n}$ of all invertible vector-polynomials (all R-automorphism of ${R_{[n]}})$ consists only of tame elements, that is, whether $GA_{n}=Tm_{n}$ or not? When $n=2$, the equality always holds; this was proved by Jung [11] for fields of characteristic zero, and was shown for positive characteristic by Van der Kulk [26]. When $n=3$ and $R$ is a field of characteristic zero, Shestakov and Umirbaiev  [21] showed that $GA_{3}\neq Tm_{3}$ by showing that the Nagata automorphism is not in $Tm_{3}$. While the case $n=3$ for positive characteristic fields and the general case $n>3$ for any field are still open. Unfortunately, the embedding of the Nagata automorphism (by adding the variables $x_{4},\ldots,x_{n}$) in the group $GA_{n}$ is shown by Smith [23] to be an element of the tame group $Tm_{n}$ for $n\geq 4$. We notice here that Maubach [14] showed that $\pi_{n}(Tm_{n})$ is just the group of all even permutations of $\mathbb{F}_{2^{r}}^{n}$ with $r>1$. He noticed that finding an element $\vec{f}\in GA_{n}$ such that $\pi_{n}(\vec{f})$ is an odd permutation is sufficient to showing $GA_{n}\neq Tm_{n}$ (see [14, Corollary 3.3]).
The case $n=1$ is much easier as in this case $Aut_{R}(R[x])=GA_{1}=Tm_{1}=\{ax+b:a,b\in R,a\neq 0\}$. However, when $R$ is not reduced, we have $Tm_{1}\neq GA_{1}$ as for example $x+rx^{2}\in GA_{1}\setminus Tm_{1}$ for every non-zero nilpotent $r$. Then may be the tameness problem is more trivial on non-reduced rings and given an $n$, there always exists a non-tame invertible element. But, we could not find a reference discussing the tameness problem on non-reduced rings. Nevertheless, this does not dampen our enthusiasm for suggesting a more general tame structure rather than the tame group. This generalized structure allows us to modify the tameness problem and to ask some related questions.
In the following, we define the tame monoid and give our general definition for tame elements.
Definition 6.1.
Let $R$ be a commutative ring with unity. We call the monoid generated by the triangular monoid $\mathcal{MT}_{n}$ and the affine group $Aff_{n}$ the tame monoid which we denote by $\mathcal{TAM}_{n}$. We call every element of $\mathcal{TAM}_{n}$ tame.
Let $\mathcal{MPP}_{n}$ denote the monoid of vector-permutation polynomials. Then, by construction, $\mathcal{MPP}_{n}$ contains $\mathcal{TAM}_{n}$. Now, we have the following general form of the tameness problem.
Question 6.2.
Let $n>1$, and let $R$ be a commutative ring with unity. Is it true that $\mathcal{MPP}_{n}=\mathcal{TAM}_{n}$?
Remark 6.3.
  1. (1)
    If $n=1$, then we always have $\mathcal{MPP}_{1}=\mathcal{TAM}_{1}$, as in this case $\mathcal{MPP}_{1}=\mathcal{MP}(R)=\mathcal{MT}_{1}$ and $Aff_{1}\subseteq\mathcal{MT}_{1}$.
  2. (2)
    If $R$ is an algebraically closed field of characteristic zero, by Proposition 4.12, we have $\mathcal{MT}_{n}=KTR_{n}$, and hence $\mathcal{MPP}_{n}=GA_{n}$. But, then Question 6.2 is reduced to the classical problem of the tame group.
We end our paper with two related questions that can be more sensible in the case of finite fields or non-reduced rings.
Question 6.4.
Let $n>1$ and let $\vec{f}\in\mathcal{MPP}_{n}$. Then does $\pi_{n}(\vec{f})\in\pi_{n}(\mathcal{TAM}_{n})$ implies that $\vec{f}\in\mathcal{TAM}_{n}$?
Another result of Maubach [14] over the finite field $\mathbb{F}_{q}$ with $q=2$ or $q$ is odd, shows that $\pi_{n}(Tm_{n})$ is the symmetric group on the elements of $\mathbb{F}_{q}^{n}$ ($n>1$). In this case, we at least know that the answer to the following question is true.
Question 6.5.
Let $n>1$. Is every vector-polynomial permutation tamely represented? In terms of symbols: Let $\vec{F}\in\pi_{n}(\mathcal{MPP}_{n})$. Then does there exist $\vec{f}\in\mathcal{TAM}_{n}$ such that $\pi_{n}(\vec{f})=\vec{F}$?
Acknowledgment. This work is supported by the Austrian Science Fund (FWF):P 35788-N. I want to express my sincere thanks to the referee for the numerous comments and valuable suggestions that improved the view of this article. Also, I would like to thank Amran A. Al-Aghbari for proof reading.